### v1

# Baltimore Atlas: FreqWeaver Adapter for Semi-supervised Ultra-high Spatial Resolution Land Cover Classification

Abstract.
Ultra-high Spatial Resolution Land Cover Classification is essential for fine-grained land cover analysis, yet it remains challenging due to the high cost of pixel-level annotations, significant scale variation, and the limited adaptability of large-scale vision models. Existing methods typically focus on 1-meter spatial resolution imagery and rely heavily on annotated data, whereas practical applications often require processing higher-resolution imagery under weak supervision. To address this, we propose a parameter-efficient semi-supervised segmentation framework for 0.3 m spatial resolution imagery, which leverages the knowledge of SAM2 and introduces a remote sensing–specific FreqWeaver Adapter to enhance fine-grained detail modeling while maintaining a lightweight design at only 5.96% of the total model parameters. By effectively leveraging unlabeled data and maintaining minimal parameter overhead, the proposed method delivers robust segmentation results with superior structural consistency, achieving a 1.78% improvement over existing parameter-efficient tuning strategies and a 3.44% gain compared to state-of-the-art high-resolution remote sensing segmentation approaches.
High-resolution remote sensing, Semi-supervised segmentation, Large model priors, Small object segmentation, Structural consistency
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: the 33rd ACM SIGSPATIAL International Conference on Advances in Geographic Information Systems; November 3–6, 2025; Minneapolis, MN, USAisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Computer visionccs: Computing methodologies Image segmentationccs: Computing methodologies Semi-supervised learningccs: Computing methodologies Transfer learningccs: Computing methodologies Convolutional neural networks
1. Introduction
High-precision segmentation of remote sensing (RS) imagery serves as a cornerstone in a wide range of real-world applications, including land cover classification, urban planning, environmental monitoring, and disaster response (Lv et al., 2023; Guo et al., 2024). With the advent of ultra-high-resolution (UHR) imagery—particularly at the 0.3-meter ground sampling distance—RS data now enables significantly richer spatial detail, allowing for the identification of fine-scale urban features such as vehicles, street furniture, residential swimming pools, and narrow alleyways (Gui et al., 2024). This spatial granularity facilitates parcel-level land use analysis and fine-grained urban modeling, thereby unlocking new frontiers for RS research and practical applications (Li et al., 2024).
However, the transition to UHR imagery introduces substantial challenges for semantic segmentation models. While existing approaches demonstrate robust performance on datasets with 1-meter resolution, their effectiveness often deteriorates when applied to finer resolutions such as 0.3 meters (Mendieta et al., 2024; Ma et al., 2025; Zhang et al., 2022). The increased level of detail exacerbates the complexity of capturing fine-grained structural boundaries and contextually diverse object classes (Huang et al., 2025). Effective segmentation at this scale requires models capable of simultaneously modeling local geometry and global semantics, motivating the development of novel architectures tailored to the intricacies of UHR data.
Moreover, a major bottleneck in the advancement of ultra-high-resolution (UHR) RS segmentation lies in the reliance on large-scale, pixel-wise annotated datasets for fully supervised training (Liu and Zhang, 2025). While effective under ideal settings, manual annotation of ultra-high-resolution RS imagery is prohibitively time-consuming and labor-intensive, limiting the scalability and practical deployment of such models (He et al., 2024). Moreover, when operating under limited supervision, these models often exhibit performance degradation and poor generalization across diverse scenes and geographies (Zhang et al., 2024).
Recently, the emergence of pre-trained foundation models such as the Segment Anything Model (SAM) (Kirillov et al., 2023a; Ravi and Gabeur, 2024) has demonstrated strong generalization capabilities across diverse visual tasks. Trained on extensive image-mask pairs, SAM2 (Ravi and Gabeur, 2024) encapsulate transferable semantic priors and offer a promising path toward alleviating annotation dependency via transfer learning on downstream tasks like high-resolution RS image segmentation (Zhang et al., 2025).
Although full fine-tuning of foundation models can enhance performance, it also incurs substantial computational and memory overheads (Han et al., 2024). Parameter-Efficient Fine-Tuning (PEFT) methods, such as Adapters (Houlsby and Giurgiu, 2019) and Low-Rank Adaptation (LoRA) (Hu and Shen, 2022), have emerged as viable alternatives for resource-conscious adaptation. However, due to the pronounced domain gap between the natural or medical images used for pre-training and the unique characteristics of RS data, conventional PEFT strategies often fail to generalize well in RS scenarios (Khanna and Irgau, 2024; Stojnic and Klemmer, 2025). Moreover, remote sensing imagery encompasses a wide spectrum of object scales—from large-scale vegetated regions to fine-grained human-made structures—necessitating robust multi-scale feature modeling capabilities (Cheng and Lei, 2023).
To this end, we propose FreqWeaver, a frequency-temporal modulation adapter designed to enhance the adaptability of foundation models like SAM2 to ultra-high-resolution (UHR) remote sensing imagery. FreqWeaver decomposes input features into low- and high-frequency components via spectral filtering, which are then processed using large and small receptive field convolutions to respectively capture global semantics and fine-grained structures. A spatial-domain residual branch is further incorporated to complement missing local details. This design not only leverages the semantic priors encoded in SAM2 for basic object recognition, but also enhances the model’s capacity to represent diverse spatial patterns in UHR data. In particular, the proposed multi-scale feature interaction mechanism significantly improves the model’s ability to delineate fine boundary details of large-scale objects and accurately identify small, morphologically complex structures—thereby enabling precise and robust segmentation across heterogeneous geographic scenes under strict parameter constraints.
Remote sensing (RS) systems, owing to their extensive spatial coverage and high temporal resolution, continuously collect vast amounts of unlabeled imagery, offering a rich yet underutilized resource for semi-supervised learning (SSL) (Yan and Yang, 2024). To address the challenges posed by limited supervision and high annotation costs in high-resolution RS segmentation, we propose a semi-supervised fine-tuning framework built upon a pre-trained foundation model. The framework integrates a teacher–student consistency learning paradigm with an uncertainty-aware weighting strategy, enabling efficient and robust domain adaptation. By dynamically modulating the contribution of pseudo-labels during training, the method effectively suppresses the adverse impact of noisy supervision—an issue particularly prevalent in RS imagery due to complex backgrounds, ambiguous boundaries, and high inter-class similarity. As a result, it significantly enhances model robustness and generalization across diverse geographic regions, acquisition conditions, and sensor types (Fan and Zhou, 2025), while simultaneously reducing reliance on costly manual annotations.
The main contributions of this paper are summarized as follows:
  • PEFT-Driven Semi-Supervised Tuning of SAM2: We propose an Adapter-based PEFT strategy integrated with an uncertainty-guided consistency regularization mechanism for the SAM2 model (Ravi and Gabeur, 2024). This approach significantly reduces the dependency on labeled data, achieving efficient segmentation performance while updating fewer than $5.96\%$ of the total model parameters.
  • 0.3m-Resolution Remote Sensing Dataset for Fine-Grained Segmentation: To support detailed land cover analysis, we construct a high-resolution (0.3m) remote sensing dataset over Baltimore. Compared to conventional 1m-Resolution datasets, it offers finer spatial granularity and greater object diversity, providing a more challenging and representative benchmark for evaluating segmentation models under real-world urban conditions.
  • FreqWeaver Adapter for Disentangled Frequency–Temporal Multi-Scale Modeling: To better capture the fine-grained structures and scale-diverse patterns in high-resolution RS imagery, we introduce FreqWeaver, a frequency–temporal modulation adapter that decomposes features into low- and high-frequency components for separate processing via large and small receptive field convolutions. A complementary spatial-domain branch further enriches semantic representation. This design enhances SAM2’s capacity to model both global context and local detail, enabling more precise and parameter-efficient multi-scale segmentation.
  • Uncertainty-Aware Semi-Supervised Learning: Considering the high error rates of pseudo-labels in remote sensing imagery—caused by complex backgrounds, fine-grained boundaries, and inter-class spectral similarity—we propose an uncertainty-aware semi-supervised framework that downweights unreliable predictions based on entropy. This design effectively suppresses misleading gradients from noisy supervision and improves the robustness of segmentation across diverse regions, sensors, and seasonal conditions.
2. Relate Work
2.1. Foundation Models in Remote Sensing
Foundation Models (FMs), notably Vision Transformer (ViT)-based architectures (Dosovitskiy et al., 2021), learn transferable representations from extensive unlabeled data, enabling strong generalization across diverse tasks. Adapting FMs to remote sensing (RS) is challenging due to RS imagery’s distinct characteristics (multi-scale spatial structures, rich spectral content, complex semantic distributions) that diverge significantly from natural images.
To address these challenges, Geospatial Foundation Models (GeoFMs) have emerged. Models like Clay (Model, 2024) and Prithvi (Szwarcman et al., 2025), pre-trained on extensive satellite/aerial data, are optimized for Earth Observation (EO). Recent advances such as S2MAE (spatial-spectral pretraining) (Placeholder et al., 2024b), GeoChat (vision-language grounding) (Placeholder et al., 2024a), and SkySense (multi-modal integration) (Placeholder et al., 2024c) highlight the trend towards task-adaptive geospatial architectures.
General-purpose FMs like the Segment Anything Model (SAM) (Kirillov et al., 2023b) and SAM2 (Ravi and Gabeur, 2024) offer prompt-driven, class-agnostic segmentation. However, the original SAM performs suboptimally on RS imagery, showing limited semantic awareness, difficulty with small/dense objects, and inconsistency in tasks like semantic change detection. These issues stem from its natural image bias and lack of RS domain priors.
To address these SAM limitations, STSAMNet (Yang et al., 2025) integrates unsupervised domain adaptation (UDA), pseudo-label refinement, and adapter layers for improved building extraction. SCNet (Jiang et al., 2025) uses multi-scale feature fusion and semantic alignment to adapt SAM2 for semantic change detection. Prompt generation and CLIP-based supervision are also explored for semantic guidance in RS.
These adaptations show a shift from universal zero-shot inference to domain-specialized fine-tuning. Thus, RS is not just adopting FMs but also shaping their development. This is driven by RS challenges like scale variation, semantic ambiguity, and the need for label-efficient geospatial analysis.
2.2. Semantic Segmentation for Complex Remote Sensing
Remote sensing image segmentation faces unique challenges. Pixel-level annotation scarcity is primary, severely limiting fully supervised methods. Second, inherent multi-scale spatial heterogeneity necessitates models adept at handling drastic scale variations (small objects to vast regions). Furthermore, high inter-class similarity and significant intra-class diversity (from varying illumination, seasons, and sensors) severely test model feature discriminability.
Semi-Supervised Semantic Segmentation (SSSS) addresses annotation scarcity by leveraging abundant unlabeled data. Key SSSS strategies include consistency regularization and pseudo-labeling(Tarvainen and Valpola, 2018; Chen et al., 2021; Sohn and Berthelot, 2020). For instance, MUCA (Wang et al., 2025) uses a Multi-scale Uncertainty Consistency (MSUC) module and cross-teacher-student attention to address multi-scale objects and inter-class similarity, enhancing multi-layer semantic consistency and discriminability. SegKC (Than et al., 2025) improves pseudo-label quality and efficiency via heterogeneous dual-network co-training with bi-directional knowledge transfer, using a lightweight network for efficient inference.
Unsupervised Domain Adaptation (UDA) addresses domain shifts in remote sensing data, often from varying sensors or locations. For example, STSAMNet (Yang et al., 2025) combines UDA with the Segment Anything Model (SAM) via an adapter, optimizing segmentation for target-domain high-resolution imagery and mitigating performance drops in new regions.
While SSSS and UDA methods partially address label scarcity and domain shifts, accurately segmenting complex spatial structures and multi-scale targets remains a bottleneck. Many methods struggle to capture diverse multi-scale features or introduce artifacts during pseudo-labeling and domain adaptation, especially at boundaries or in textured areas. To overcome these limitations, we explore parameter-efficient solutions that not only reduce reliance on labeled data but also enhance multi-scale representation under minimal supervision, as detailed next.
2.3. Parameter-Efficient Fine-Tuning Method
As foundation models (FMs), including geospatial variants (GeoFMs), continue to grow in size and complexity, full fine-tuning has become increasingly impractical due to high computational costs and risks such as overfitting and catastrophic forgetting. Parameter-Efficient Fine-Tuning (PEFT) offers a scalable alternative by adapting models through the selective update of a small subset of parameters, while keeping the majority of pre-trained weights frozen.
Several PEFT techniques have been proposed to balance efficiency and performance. Adapter tuning, which inserts lightweight modules into transformer layers, has been successfully applied in models like STSAMNet (Yang et al., 2025) for adapting SAM to high-resolution remote sensing. Low-Rank Adaptation (LoRA) (Hu et al., 2021) reduces parameter overhead by constraining updates to low-rank subspaces. Prompt tuning (Jia et al., 2022) and prefix tuning (Li and Liang, 2021) modify task-specific behavior through learned embeddings, without altering the base model. These methods significantly reduce training cost and memory usage, and often enhance generalization by preserving pre-trained representations.
PEFT has gained traction in remote sensing. (Marti-Escofet et al., 2025) benchmarked multiple PEFT methods across diverse GeoFM architectures and Earth observation datasets, showing that PEFT can match or surpass full fine-tuning with far lower resource requirements. (He et al., 2023) introduced Sensitivity-aware PEFT (SPT), which identifies and prioritizes task-relevant parameters for adaptation, integrating LoRA or adapters accordingly. Domain-specific approaches such as AiRs (Hu et al., 2024b), TEA (Hu et al., 2024a), and Earth-Adapter (Hu et al., 2025) further enhance transformer adaptation for Earth observation via residual-enhanced adapters.
To address key challenges in remote sensing imagery—such as large parameter overhead, aliasing artifacts, and significant object-scale variability—we propose a novel adapter architecture that combines frequency-domain decomposition with specialized convolutional processing. By disentangling high- and low-frequency components through dedicated convolutional branches and incorporating temporal information, the proposed method enables effective multi-scale adaptation while maintaining minimal parameter cost.
Figure 1. Overview of the proposed framework.
3. Method
We propose a semi-supervised, parameter-efficient fine-tuning framework built upon the Segment Anything Model 2 (SAM2), which effectively leverages the prior knowledge of large-scale foundation models. To this end, we introduce the FreqWeaver Adapter—a lightweight yet effective module that adapts SAM2 to remote sensing data while maintaining high parameter efficiency. In addition, we design a semi-supervised learning strategy that fully utilizes unlabeled data to improve domain adaptation and generalization under limited supervision. The overall framework is illustrated in Figure 1.
3.1. FreqWeaver: Frequency-Modulated Adapter for Efficient RS Fine-Tuning
Remote sensing (RS) imagery is typically characterized by significant scale variation, densely distributed objects, and complex textures. These properties pose challenges not only to general-purpose vision models such as SAM, but also to existing parameter-efficient tuning methods, which often struggle to adapt to such complex data distributions. To address this, we propose FreqWeaver, a lightweight frequency–temporal modulation adapter. FreqWeaver explicitly decomposes input features into low- and high-frequency components, which are processed separately using convolutional filters with distinct receptive fields to model global semantics and fine-grained details, respectively. In addition, a temporal branch is introduced to capture residual semantic information that may not be effectively represented in the frequency domain. This design significantly improves the adaptability and structural representation capability of SAM on RS imagery, while maintaining parameter efficiency.An overview of the proposed adapter is illustrated in Figure 2.
3.1.1. Spectral Decomposition
To facilitate frequency-aware feature refinement, we perform spectral decomposition on the encoder outputs using the 2D Discrete Fourier Transform (DFT). This allows explicit separation of low- and high-frequency information for targeted processing.
Before applying the DFT, we first reshape the input feature tensor from $\mathbf{X}\in\mathbb{R}^{B\times H\times W\times C}$ to $\mathbf{X}^{\prime}\in\mathbb{R}^{B\times C\times H\times W}$ to align with the channel-wise Fourier transform convention.
We then apply the 2D DFT independently to each channel as:
(1) $\mathbf{X}_{\mathrm{f}}=\mathcal{F}_{\mathrm{DFT}}(\mathbf{X}^{\prime}),$
where $\mathcal{F}_{\mathrm{DFT}}$ denotes the 2D Fourier transform along the spatial dimensions $(H,W)$ for each channel. The resulting spectrum $\mathbf{X}_{\mathrm{f}}\in\mathbb{C}^{B\times C\times H\times W}$ contains complex-valued frequency coefficients.
To extract different frequency bands, we define two complementary radial binary masks:
(2) $\mathcal{M}_{\mathrm{lf}}=\mathbb{I}(r<\rho),\quad\mathcal{M}_{\mathrm{hf}}=1-\mathcal{M}_{\mathrm{lf}},$
where $r$ denotes the radial distance from the spectral center, and $\rho$ is a predefined cutoff threshold.
We isolate the low- and high-frequency components by applying the masks element-wise:
(3) $\mathbf{X}_{\mathrm{lf}}=\mathbf{X}_{\mathrm{f}}\odot\mathcal{M}_{\mathrm{lf}},\quad\mathbf{X}_{\mathrm{hf}}=\mathbf{X}_{\mathrm{f}}\odot\mathcal{M}_{\mathrm{hf}},$
where $\odot$ denotes Hadamard (element-wise) multiplication.
Finally, we recover the spatial-domain representations via the inverse Fourier transform:
(4) $\widehat{\mathbf{X}}_{\mathrm{lf}}=\mathcal{F}^{-1}_{\mathrm{DFT}}(\mathbf{X}_{\mathrm{lf}}),\quad\widehat{\mathbf{X}}_{\mathrm{hf}}=\mathcal{F}^{-1}_{\mathrm{DFT}}(\mathbf{X}_{\mathrm{hf}}),$
resulting in real-valued tensors $\widehat{\mathbf{X}}_{\mathrm{lf}},\widehat{\mathbf{X}}_{\mathrm{hf}}\in\mathbb{R}^{B\times C\times H\times W}$ representing the low- and high-frequency spatial features, respectively. These components are subsequently passed to dedicated branches for frequency-specific refinement.
Figure 2. Overview of the proposed FreqWeaver Adapter.
3.1.2. Frequency–Temporal Feature Modulation
To enhance frequency- and spatial-domain representations in a disentangled manner, FreqWeaver introduces three specialized branches, each responsible for modeling a specific component of the feature spectrum: low-frequency (global structure), high-frequency (local detail), and spatial-domain semantics. These branches operate independently and produce complementary signals for subsequent fusion.
Low-Frequency Branch.
This branch targets the restoration of global semantic coherence and suppression of large-scale distortions, which are commonly observed in low-frequency-dominated regions such as terrain or vegetation. By modeling long-range dependencies, it reinforces region-level consistency across the scene.
The low-frequency response is computed as:
(5) $\mathbf{F}_{\mathrm{lf}}=\mathcal{F}_{\mathrm{LF}}(\widehat{\mathbf{X}}_{\mathrm{lf}}),$
where $\widehat{\mathbf{X}}_{\mathrm{lf}}\in\mathbb{R}^{B\times C\times H\times W}$ is the input low-frequency feature map, and $\mathcal{F}_{\mathrm{LF}}(\cdot)$ denotes a depthwise convolution with a kernel size of $11\times 1$ 11, initialized with zeros. The output $\mathbf{F}_{\mathrm{lf}}\in\mathbb{R}^{B\times C\times H\times W}$ encodes coarse structural information and is reshaped to $\mathbb{R}^{B\times H\times W\times C}$.
High-Frequency Branch.
This branch emphasizes fine-grained details such as edges, contours, and textures—high-frequency components that are essential for distinguishing man-made objects and detailed structures. It enhances boundary clarity and preserves local precision.
The high-frequency response is defined as:
(6) $\mathbf{F}_{\mathrm{hf}}=\mathcal{F}_{\mathrm{HF}}(\widehat{\mathbf{X}}_{\mathrm{hf}}),$
where $\widehat{\mathbf{X}}_{\mathrm{hf}}\in\mathbb{R}^{B\times C\times H\times W}$ is the high-frequency input, and $\mathcal{F}_{\mathrm{HF}}(\cdot)$ is a depthwise convolution with a $5\times 5$ kernel. The output $\mathbf{F}_{\mathrm{hf}}\in\mathbb{R}^{B\times C\times H\times W}$ highlights spatially localized structures and is similarly reshaped to $\mathbb{R}^{B\times H\times W\times C}$.
Spatial Branch.
This branch directly processes the spatial-domain feature without spectral filtering, enabling the modeling of mid-frequency signals and ambiguous semantic cues that may not be cleanly separated in the frequency domain. It plays a complementary role in capturing transitional or occluded content.
The spatial-domain response is formulated as:
(7) $\mathbf{F}_{\mathrm{s}}=\mathcal{F}_{\mathrm{S}}(\mathbf{X}),$
where $\mathbf{X}\in\mathbb{R}^{B\times H\times W\times C}$ is the original feature map before frequency decomposition. The operator $\mathcal{F}_{\mathrm{S}}(\cdot)$ is a two-layer bottleneck MLP. The output $\mathbf{F}_{\mathrm{s}}\in\mathbb{R}^{B\times H\times W\times C}$ captures spatial-domain semantics that complement the frequency-based branches.
3.1.3. Content-Aware Feature Fusion
To integrate the outputs of the frequency–temporal branches, we adopt a content-aware fusion strategy guided by global contextual information. This mechanism allows the model to dynamically balance the relative contributions of low-frequency, high-frequency, and spatial-domain features on a per-sample basis.
We first compute a global context descriptor via global average pooling over the spatial dimensions:
(8) $\mathbf{g}=\frac{1}{HW}\sum_{h=1}^{H}\sum_{w=1}^{W}\mathbf{X}_{:,h,w,:}\in\mathbb{R}^{B\times C},$
where $\mathbf{X}\in\mathbb{R}^{B\times H\times W\times C}$ is the original spatial-domain feature map.
The descriptor $\mathbf{g}$ is then passed to a lightweight router MLP, which outputs normalized fusion weights:
(9) $[r_{\mathrm{s}},r_{\mathrm{lf}},r_{\mathrm{hf}}]=\operatorname{softmax}(\mathcal{F}_{\mathrm{Router}}(\mathbf{g}))\in\mathbb{R}^{B\times 3},$
where $\mathcal{F}_{\mathrm{Router}}(\cdot)$ denotes a two-layer MLP with softmax normalization to ensure $\sum_{k}r_{k}=1$.
Finally, the three modulation outputs are adaptively fused with the original features via weighted addition:
(10) $\mathbf{Y}=\mathbf{X}+r_{\mathrm{s}}\odot\mathbf{F}_{\mathrm{s}}+r_{\mathrm{lf}}\odot\mathbf{F}_{\mathrm{lf}}+r_{\mathrm{hf}}\odot\mathbf{F}_{\mathrm{hf}},$
where $\odot$ denotes broadcasted element-wise multiplication. The final output $\mathbf{Y}\in\mathbb{R}^{B\times H\times W\times C}$ preserves the original spatial resolution while incorporating frequency- and content-aware enhancements.
This fusion strategy enables adaptive emphasis of complementary information sources, thereby improving the model’s ability to generalize to diverse spatial patterns in remote sensing imagery.
3.2. Uncertainty-Aware Teacher–Student Adaptation
In the context of remote sensing segmentation, acquiring large-scale annotated datasets is both expensive and labor-intensive. To alleviate this constraint, we introduce an Uncertainty-Aware Mean Teacher (UAMT) framework that leverages unlabeled data via a dual-network consistency paradigm. The method enhances pseudo-supervision reliability by integrating entropy-based uncertainty estimation into a perturbation-consistent learning scheme. This design allows for stable adaptation of large vision foundation models (e.g., SAM2) with minimal supervision.
3.2.1. Mean Teacher Framework with Perturbation Consistency
Our framework adopts a teacher–student architecture where the student network $\boldsymbol{\theta}_{\mathrm{S}}$ is optimized via gradient descent, while the teacher network $\boldsymbol{\theta}_{\mathrm{T}}$ acts as a temporal ensemble, updated using Exponential Moving Average (EMA) to ensure stable supervision:
(11) $\boldsymbol{\theta}_{\mathrm{T}}^{(t)}=\alpha\,\boldsymbol{\theta}_{\mathrm{T}}^{(t-1)}+(1-\alpha)\,\boldsymbol{\theta}_{\mathrm{S}}^{(t)},\quad\alpha=0.99,$
where $\alpha$ is the EMA decay coefficient controlling the momentum of the update.
To enforce robustness under data augmentation, each unlabeled input $x_{u}\in\mathcal{D}_{u}$ is subjected to two augmentation pipelines: a weakly augmented version $\widetilde{x}^{\text{weak}}_{u}$ is fed to the teacher model, while a strongly augmented counterpart $\widetilde{x}^{\text{strong}}_{u}$ is provided to the student. The underlying assumption is that semantic predictions should remain consistent across these perturbations, even if the appearance varies. Accordingly, consistency regularization is applied to align the outputs of teacher and student networks under this augmentation mismatch.
3.2.2. Entropy-Based Uncertainty Estimation
While teacher predictions provide valuable supervisory cues, pseudo-labels generated for remote sensing images are particularly error-prone due to complex backgrounds, fine-grained object boundaries, and domain-specific variations. To mitigate the impact of such noisy supervision, we employ a pixel-wise confidence estimation mechanism based on Shannon entropy to downweight uncertain predictions.
Let $\mathbf{P}^{\text{T}}_{j}\in\mathbb{R}^{C}$ denote the softmax probability vector predicted by the teacher at pixel $j$, where $C$ is the number of classes. The uncertainty score is quantified by entropy:
(12) $\mathcal{H}_{j}=-\sum_{c=1}^{C}\mathbf{P}^{\text{T}}_{j,c}\log\mathbf{P}^{\text{T}}_{j,c},$
which attains its maximum at $\log C$ when the prediction is maximally ambiguous (i.e., uniform distribution). To normalize this measure and convert it into a usable confidence weight, we define:
(13) $w_{j}=1-\frac{\mathcal{H}_{j}}{\log C},\quad w_{j}\in[0,1],$
where higher $w_{j}$ values indicate more confident predictions and hence stronger training signals. This design enables the model to suppress the impact of high-entropy (uncertain) regions and focus on reliable pseudo-labels.
3.2.3. Loss Function: Supervised–Unsupervised Coupling
The overall objective consists of two complementary terms: a supervised loss on labeled data and a confidence-weighted consistency loss on unlabeled data.
Supervised Loss.
Given labeled samples $(x_{l},y_{l})\in\mathcal{D}_{l}$, the student model is trained with standard pixel-wise cross-entropy:
(14) $\mathcal{L}_{\mathrm{sup}}=\mathrm{CE}\left(f(x_{l};\boldsymbol{\theta}_{\mathrm{S}}),y_{l}\right),$
where $f(\cdot;\boldsymbol{\theta}_{\mathrm{S}})$ denotes the student prediction.
Consistency Loss with Confidence Weighting.
For unlabeled images $x_{u}\in\mathcal{D}_{u}$, the teacher generates soft predictions $\mathbf{P}^{\text{T}}$, while the student predicts $\mathbf{P}^{\text{S}}$ from a strongly perturbed view. A pseudo-label $\hat{y}^{\text{T}}_{j}=\arg\max_{c}\mathbf{P}^{\text{T}}_{j,c}$ is assigned per pixel, and the confidence-weighted consistency loss is formulated as:
(15) $\mathcal{L}_{\mathrm{cons}}=\frac{1}{|\mathcal{D}_{u}|\cdot N}\sum_{x_{u}\in\mathcal{D}_{u}}\sum_{j=1}^{N}w_{j}\cdot\mathrm{CE}(\mathbf{P}^{\text{S}}_{j},\hat{y}^{\text{T}}_{j}),$
where $N$ is the total number of pixels per image, and $\mathbf{P}^{\text{S}}_{j}$ is the predicted softmax from the student at pixel $j$.
Total Objective.
The combined loss is given by:
(16) $\mathcal{L}_{\mathrm{total}}=\mathcal{L}_{\mathrm{sup}}+\lambda\cdot\mathcal{L}_{\mathrm{cons}},$
where $\lambda=0.1$ balances the contributions of labeled and unlabeled samples.
During training, only the parameters of the FreqWeaver adapters are updated, while the SAM2 backbone remains frozen. This strategy ensures efficient and stable transfer to remote sensing domains, leveraging both labeled and unlabeled data in a principled manner.
Figure 3. Visualization of the State-of-the-Art PEFT Method.
4. Experiment
4.1. Dataset
In this study, all experiments were conducted on ultra-high-resolution remote sensing imagery of Baltimore, captured at a spatial resolution of 0.3 meters. A total of 60 images, each with dimensions of 1024×1024 pixels, were selected based on the nine urban semantic categories defined by UrbanWatch, ensuring a relatively balanced distribution of class labels, as shown in Fig 4. The dataset comprises a total of 62,914,560 pixels.For initial annotations, we utilized the UrbanWatch model, which was originally trained on 1-meter resolution imagery. Although not optimized for ultra-high-resolution inputs, this model offered a coarse yet informative baseline. To improve annotation accuracy, we subsequently performed meticulous manual corrections. Given the complexity of remote sensing scenes and the difficulty of pixel-level labeling, this process was both labor-intensive and time-consuming.For training and evaluation, we used 55 images for training and the remaining 5 for testing. The test set alone contained 5,242,880 pixels, offering a sufficiently large sample size for robust and reliable performance evaluation.
Figure 4. Visualization of land cover class distribution.
4.2. Evaluation Metrics
To comprehensively evaluate the segmentation performance, we adopt two widely used metrics: Intersection over Union (IoU) and Dice coefficient. Both metrics quantify the region-level overlap between predicted segmentation results and ground truth annotations, providing a general assessment of segmentation accuracy.
Given the predicted region $P$ and the ground truth region $G$, the IoU and Dice coefficient are defined as follows:
(17) $\text{IoU}=\frac{|P\cap G|}{|P\cup G|},$
(18) $\text{Dice}=\frac{2|P\cap G|}{|P|+|G|},$
where $|P\cap G|$ denotes the number of pixels in the intersection of the prediction and the ground truth, and $|P\cup G|$ denotes the number of pixels in their union. Higher values of IoU and Dice indicate better segmentation performance.
4.3. Training and Testing Details
All experiments were conducted using PyTorch on a single NVIDIA RTX L40S GPU with 48 GB of memory. In our work, we adopt SAM2 as the backbone network. Input images were uniformly resized to $1024\times 1024$ during both training and testing. The model was trained for 200 epochs using the Adam optimizer with a learning rate of $5\times 10^{-4}$ and a batch size of 2.
For the segmentation task, we employed the standard Cross-Entropy loss, placing emphasis on accurate pixel-wise classification as the primary optimization objective.
Table 1. Comparison with state-of-the-art parameter-efficient fine-tuning (PEFT) methods.
Methods IoU Dice Param
Full Fine-tuning 0.5239 0.6006 69.11M
Earth Adapter (2025) (Hu et al., 2025) 0.5149 0.5893 10.19M
LoRA (2021) (Hu and Shen, 2022) 0.5202 0.5909 0.48M
SAM Adapter (2023) (Wu et al., 2023b) 0.5321 0.6116 5.64M
Conv-LoRA (2024) (Zhong et al., 2024) 0.5342 0.6051 0.50M
Ours 0.5520 0.6331 4.12M
Figure 5. Visualization of the State-of-the-Art LULC Method.
4.4. Comparison With State-of-the-Art PEFT Methods
To evaluate the effectiveness of the proposed adapter, we conducted extensive comparative experiments against several state-of-the-art parameter-efficient fine-tuning (PEFT) methods, including the SAM Adapter(Wu et al., 2023b), LoRA(Hu and Shen, 2022), Conv-LoRA(Zhong et al., 2024), and the Earth Adapter(Hu et al., 2025). These approaches represent diverse strategies for adapting large-scale models like SAM2, with different focuses on domain specificity and parameter efficiency, as summarized in Table 1.
Limitations of LoRA and Conv-LoRA. Our experimental results show that both LoRA and Conv-LoRA achieve competitive performance with remarkably low parameter overhead. LoRA, which leverages low-rank decomposition for efficient adaptation, is widely adopted for lightweight fine-tuning tasks. However, due to its linear structure, it exhibits limited capability in modeling the spatial complexity of high-resolution remote sensing scenes, which often involve multi-scale objects, intricate semantics, and fine-grained textures. To address this limitation, Conv-LoRA enhances the original LoRA design by incorporating convolutional layers to strengthen local feature extraction. This modification improves its ability to model small objects and boundary details, leading to noticeable performance gains. Nonetheless, despite these improvements, Conv-LoRA still lacks the representational capacity to fully exploit the potential of high-resolution remote sensing imagery, particularly in capturing complex spatial-frequency interactions, leaving room for further enhancement.
Limitations of Earth Adapter and SAM Adapter. The Earth Adapter, although specifically designed for remote sensing tasks, attempts to decouple frequency components without incorporating targeted spatial modeling mechanisms. This results in relatively high parameter consumption and unsatisfactory performance. The SAM Adapter, originally developed for medical image segmentation, shows moderate adaptability when transferred to remote sensing domains. However, due to the increased complexity and resolution of remote sensing data, it struggles to effectively model fine-grained structures and semantic variations.
Effectiveness of the Proposed FreqWeaver Adapter. In contrast, our proposed FreqWeaver Adapter introduces a frequency-aware design tailored for remote sensing applications. By explicitly decoupling low- and high-frequency components and applying targeted processing strategies to each, FreqWeaver effectively suppresses low-frequency artifacts while preserving high-frequency structural details. This leads to improved segmentation accuracy and better adaptation to the SAM2 backbone. Importantly, FreqWeaver achieves this superior performance with fewer parameters than both the Earth Adapter and the SAM Adapter, demonstrating its strong domain-specific effectiveness and efficiency for high-resolution remote sensing tasks. We also present visualizations to validate the effectiveness of our method, as shown in 3.
Table 2. Comparison with state-of-the-art Land Use and Land Cover (LULC) methods.
Methods IoU Dice
ABCNet(2021)(Li et al., 2021) 0.4111 0.5217
UNetFormer(2022)(Wang et al., 2022) 0.5236 0.6201
CMTFNet(2023)(Wu et al., 2023a) 0.5428 0.6369
RS3Mamba(2024)(Ma et al., 2024b) 0.5092 0.6068
SAM-RS(2024)(Ma et al., 2024a) 0.5434 0.6342
Ours 0.5778 0.6545
4.5. Comparison with State-of-the-Art Land Cover Classification Methods
To evaluate the effectiveness of our proposed approach, we conducted extensive comparative experiments against several representative state-of-the-art Land Cover and Land Use Classification (LULC) methods. Specifically, we selected five recent and competitive baselines from the literature (2021–2024)—ABCNet (Li et al., 2021), UNetFormer (Wang et al., 2022), CMTFNet (Wu et al., 2023a), RS3Mamba (Ma et al., 2024b), and SAM-RS (Ma et al., 2024a) chosen for their strong benchmark performance and representation of diverse architectural paradigms, including CNN-based, Transformer-based, and Mamba-based models. The quantitative comparison results are reported in Table 2, while qualitative visualization examples are provided in Fig. 5 to further illustrate the advantages of our method.
Limitations of Existing Methods. Quantitative results (Table 2) demonstrate that the proposed method consistently outperforms all selected baseline methods. In comparison to the strongest baseline, SAM-RS, our model achieves an improvement of 3.44% in IoU and 2.03% in Dice score, underscoring its superior performance in LULC tasks. Despite the reasonable segmentation accuracy exhibited by prior methods, visual analysis reveals several limitations: (1) Misclassification remains prevalent in complex scenes, particularly those with high intra-class variance or subtle inter-class distinctions (e.g., differentiating various types of vegetation or urban structures); (2) Object boundaries are often rendered blurry and imprecise, a critical deficiency for ultra-high-resolution segmentation tasks that demand sharp delineations; (3) Models generally perform well on spectrally distinct classes like grass and trees, but falter with categories exhibiting complex geometries or semantic overlaps, such as distinguishing between parking lots and buildings, or between roads and barren land. Moreover, small object detection, such as identifying pools in residential areas, is often overlooked by existing approaches.
Effectiveness of Our Method. To the best of our knowledge, our model is the first to adapt the Segment Anything v2 (SAM2) architecture for the task of LULC. By leveraging the rich prior knowledge encoded in this foundation model, we obtain robust category- and pixel-level representations. Building upon this foundation, the proposed FreqWeaver Adapter enhances the semantic granularity of feature modeling—by explicitly capturing and integrating multi-frequency information crucial for detail preservation—thereby enabling significantly improved delineation of object boundaries and detection of small structures. Notably, the enhanced feature representation from the FreqWeaver Adapter allows our model to effectively exploit contextual cues, such as the presence or absence of vehicles, to better distinguish between semantically similar regions like parking lots and roads. These capabilities demonstrate the strengths of our architecture in ultra-high-resolution LULC tasks. Additionally, the integration of a semi-supervised learning strategy further improves the robustness and generalization of our model under limited annotation scenarios.
Table 3. Classification performance on ultra-high resolution Downstream Applications.
Class User’s Accuracy (%) Producer’s Accuracy (%)
Building 93.71 81.71
Road 85.39 89.94
Parking lot 85.00 85.53
Tree Canopy 75.58 98.98
Grass/Shrub 91.80 73.36
Agriculture 0.00 0.00
Water 94.92 98.25
Barren 80.00 40.00
Other 60.00 54.55
Overall Accuracy (OA) 84.55%
Kappa Coefficient 81.38%
Figure 6. Visualization of Land Cover Classification.
4.6. Exploration of Downstream Applications in Map Analysis
To evaluate the practicality of our approach in real-world scenarios, we applied the proposed method to downstream ultra-high-resolution map analysis tasks, as summarized in Table 3. Despite being trained on a relatively small dataset, our method exhibits strong generalization capabilities, achieving an Overall Accuracy (OA) of 84.55% and a Kappa coefficient of 81.38%. These metrics reflect not only the correctness of predictions but also the statistical reliability of classification performance across diverse land cover categories.
A detailed breakdown of class-wise performance is provided in Table 3. Notably, high User’s Accuracy was achieved for categories such as Building (93.71%), Water (94.92%), and Grass/Shrub (91.80%), indicating precise identification of spatial regions with distinct spectral or structural characteristics. At the same time, balanced Producer’s Accuracy was maintained, particularly for Tree Canopy (98.98%) and Road (89.94%), demonstrating strong recall and comprehensive spatial coverage.
Nevertheless, certain challenges persist in distinguishing semantically overlapping or visually ambiguous regions, such as Tree Canopy vs. Grass/Shrub and Parking Lot vs. Road, which are frequently misclassified due to spatial proximity or similar textures.
To further validate the effectiveness of our method, qualitative visualizations were performed on real-world map segments, as illustrated in Fig. 6. The results show that our approach successfully preserves object boundaries, captures fine-grained structures, and adapts effectively to spatial heterogeneity, confirming its suitability for downstream high-resolution map analysis tasks.
Table 4. Quantitative results of the ablation study.
FW UATS IoU Dice
0.5239 0.6006
0.5520 0.6331
0.5778 0.6545
4.7. Ablation Studies
To evaluate the contribution of each component within the proposed framework, we conducted extensive ablation studies, as shown in Table 4. In each experiment, specific modules—FreqWeaver (FW) and the Uncertainty-Aware Teacher Student Model (UATS)—were selectively enabled or disabled to analyze their individual impact on segmentation performance.
Effectiveness of FreqWeaver. We introduce FreqWeaver (FW) to decouple the frequency and temporal components within SAM2, enabling a parameter-efficient mechanism to capture complex patterns in remote sensing imagery. Experimental results show that, compared to full fine-tuning, our method achieves a 2.81% improvement in IoU and a 3.25% improvement in Dice score, while updating only 5.96% of the total parameters. These findings highlight the effectiveness of FreqWeaver in fine-tuning large vision models for remote sensing tasks and provide a scalable pathway for adapting foundation models under strict resource constraints.
Effectiveness of the Uncertainty-Aware Teacher–Student Model. We introduce the Uncertainty-Aware Teacher–Student model (UATS) to mitigate the impact of unreliable pseudo-labels, which are common in remote sensing imagery due to complex backgrounds and ambiguous boundaries. By incorporating uncertainty estimation into the consistency learning framework, UATS effectively downweights noisy supervision and makes better use of unlabeled data in the semi-supervised setting. As shown in our experiments, adding UATS on top of FreqWeaver leads to a 2.58% improvement in IoU and a 2.14% improvement in Dice score. These results demonstrate the importance of uncertainty modeling in enhancing model robustness and generalization across diverse remote sensing scenes.
Table 5. Quantitative results of the Adaptation ablation study.
Stage 1 Stage 2 Stage 3 Stage 4 IoU Dice
0.5238 0.5997
0.4955 0.5694
0.5520 0.6331
0.3722 0.4633
4.8. Segment Anything 2 Adaptation
Segment Anything 2 is built upon the Hierarchical Detection Transformer (HieraDet), a multi-scale transformer architecture composed of four stages. In the original Segment Anything framework, adapters such as the SAM Adapter are inserted into every transformer block. However, directly applying this strategy to Segment Anything 2 often results in suboptimal fine-tuning performance due to its deeper and hierarchically structured design. To better accommodate these architectural differences, we conducted a series of empirical studies to investigate more effective adapter integration strategies. The results are presented in Table 5.
Specifically, we hypothesize that fine-tuning a foundation model to adapt to downstream tasks such as remote sensing segmentation primarily requires the adjustment of deeper semantic layers. Therefore, we initially inserted adapters into all four stages and gradually moved toward inserting them only in the deeper stages. Experimental results reveal that adding adapters to early transformer stages not only fails to improve domain adaptation, but can even hinder the model’s ability to specialize for remote sensing data. Conversely, limiting adapter placement to only the final stage also leads to underfitting, as the model lacks sufficient capacity for structural adaptation. We observe that inserting adapters into Stage 3 and Stage 4 yields the best performance, striking a balance between semantic depth and structural flexibility. While our experiments primarily report results using the proposed FreqWeaver Adapter, we also validated this stage-wise trend using other adapter designs, including the SAM Adapter and Earth Adapter, and found consistent behavior across all cases.
5. Conclusion
In this work, we proposed a parameter-efficient semi-supervised segmentation framework for high-resolution remote sensing imagery. By integrating a domain-adaptive FreqWeaver Adapter with the SAM2 backbone and employing an uncertainty-aware teacher student strategy, our method addresses the challenges of limited annotations, scale variation, and structural consistency. Experiments show that our approach achieves strong segmentation performance with minimal parameter overhead, outperforming existing PEFT and high-resolution segmentation methods. These results highlight the potential of combining spectral decomposition and semi-supervised learning for effective foundation model adaptation in remote sensing.
6. Discussion
One limitation of our current study lies in the geographic scope of the dataset, which is primarily extracted from selected regions within Baltimore City. Although class balancing techniques were applied during preprocessing, certain land cover categories—such as agriculture—remain significantly underrepresented. This imbalance may introduce bias into the model’s learning process, particularly in recognizing rare or sparsely distributed classes. In future work, we plan to extend our dataset to comprehensively cover the entire Baltimore City area. By doing so, we aim to achieve a more balanced and representative distribution of land cover classes, which is expected to enhance the generalizability and robustness of the segmentation model.
Acknowledgements.
This research is part of the funded project ”BaltimoreAtlas: An ultra-high-resolution land cover and land use classification for the Greater Baltimore Metropolitan Area”, supported by the School of Emerging Technologies (SET) seed funding, awarded by Towson University. Additional data acquisition, processing, and analysis were supported by the U.S. Geological Survey under Grant/Cooperative Agreement No. G23AP00683 (GY23-GY27).

### v2

@@
 Abstract.
-Ultra-high Spatial Resolution Land Cover Classification is essential for fine-grained land cover analysis, yet it remains challenging due to the high cost of pixel-level annotations, significant scale variation, and the limited adaptability of large-scale vision models.
-Existing methods typically focus on 1-meter spatial resolution imagery and rely heavily on annotated data, whereas practical applications often require processing higher-resolution imagery under weak supervision.
-To address this, we propose a parameter-efficient semi-supervised segmentation framework for 0.3 m spatial resolution imagery, which leverages the knowledge of SAM2 and introduces a remote sensing–specific FreqWeaver Adapter to enhance fine-grained detail modeling while maintaining a lightweight design at only 5.96% of the total model parameters.
-By effectively leveraging unlabeled data and maintaining minimal parameter overhead, the proposed method delivers robust segmentation results with superior structural consistency, achieving a 1.78% improvement over existing parameter-efficient tuning strategies and a 3.44% gain compared to state-of-the-art high-resolution remote sensing segmentation approaches.
-High-resolution remote sensing, Semi-supervised segmentation, Large model priors, Small object segmentation, Structural consistency
-copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: the 33rd ACM SIGSPATIAL International Conference on Advances in Geographic Information Systems; November 3–6, 2025; Minneapolis, MN, USAisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Computer visionccs: Computing methodologies Image segmentationccs: Computing methodologies Semi-supervised learningccs: Computing methodologies Transfer learningccs: Computing methodologies Convolutional neural networks
+Ultra-high Spatial Resolution (UHSR) Land Cover Classification is increasingly important for urban analysis, enabling fine-scale planning, ecological monitoring, and infrastructure management.
+It identifies land cover types on sub-meter remote sensing imagery, capturing details such as building outlines, road networks, and distinct boundaries.
+However, most existing methods focus on 1 m imagery and rely heavily on large-scale annotations, while UHSR data remain scarce and difficult to annotate, limiting practical applicability.
+To address these challenges, we introduce Baltimore Atlas, a UHSR land cover classification framework that reduces reliance on large-scale training data and delivers high-accuracy results.
+Baltimore Atlas builds on three key ideas: (1) Baltimore Atlas Dataset, a 0.3 m resolution dataset based on aerial imagery of Baltimore City; (2) FreqWeaver Adapter, a parameter-efficient adapter that transfers SAM2 to this domain, leveraging foundation model knowledge to reduce training data needs while enabling fine-grained detail and structural modeling; (3) Uncertainty-Aware Teacher Student Framework, a semi-supervised framework that exploits unlabeled data to further reduce training dependence and improve generalization across diverse scenes.
+Using only 5.96% of total model parameters, our approach achieves a 1.78% IoU improvement over existing parameter-efficient tuning strategies and a 3.44% IoU gain compared to state-of-the-art high-resolution remote sensing segmentation methods on the Baltimore Atlas Dataset.
+Ultra-high Spatial Resolution Land Cover Classification, Parameter-Efficient Adapter, Semi-Supervised Segmentation, Transfer Learning
+journalyear: 2025copyright: ccconference: The 33rd ACM International Conference on Advances in Geographic Information Systems; November 3–6, 2025; Minneapolis, MN, USAbooktitle: The 33rd ACM International Conference on Advances in Geographic Information Systems (SIGSPATIAL ’25), November 3–6, 2025, Minneapolis, MN, USAdoi: 10.1145/3748636.3762724isbn: 979-8-4007-2086-4/2025/11ccs: Computing methodologies Computer visionccs: Computing methodologies Image segmentationccs: Computing methodologies Semi-supervised learning settingsccs: Information systems Geographic information systemsccs: Theory of computation Design and analysis of algorithms
 1. Introduction
-High-precision segmentation of remote sensing (RS) imagery serves as a cornerstone in a wide range of real-world applications, including land cover classification, urban planning, environmental monitoring, and disaster response (Lv et al., 2023; Guo et al., 2024).
-With the advent of ultra-high-resolution (UHR) imagery—particularly at the 0.3-meter ground sampling distance—RS data now enables significantly richer spatial detail, allowing for the identification of fine-scale urban features such as vehicles, street furniture, residential swimming pools, and narrow alleyways (Gui et al., 2024).
-This spatial granularity facilitates parcel-level land use analysis and fine-grained urban modeling, thereby unlocking new frontiers for RS research and practical applications (Li et al., 2024).
-However, the transition to UHR imagery introduces substantial challenges for semantic segmentation models.
-While existing approaches demonstrate robust performance on datasets with 1-meter resolution, their effectiveness often deteriorates when applied to finer resolutions such as 0.3 meters (Mendieta et al., 2024; Ma et al., 2025; Zhang et al., 2022).
-The increased level of detail exacerbates the complexity of capturing fine-grained structural boundaries and contextually diverse object classes (Huang et al., 2025).
-Effective segmentation at this scale requires models capable of simultaneously modeling local geometry and global semantics, motivating the development of novel architectures tailored to the intricacies of UHR data.
-Moreover, a major bottleneck in the advancement of ultra-high-resolution (UHR) RS segmentation lies in the reliance on large-scale, pixel-wise annotated datasets for fully supervised training (Liu and Zhang, 2025).
-While effective under ideal settings, manual annotation of ultra-high-resolution RS imagery is prohibitively time-consuming and labor-intensive, limiting the scalability and practical deployment of such models (He et al., 2024).
-Moreover, when operating under limited supervision, these models often exhibit performance degradation and poor generalization across diverse scenes and geographies (Zhang et al., 2024).
-Recently, the emergence of pre-trained foundation models such as the Segment Anything Model (SAM) (Kirillov et al., 2023a; Ravi and Gabeur, 2024) has demonstrated strong generalization capabilities across diverse visual tasks.
-Trained on extensive image-mask pairs, SAM2 (Ravi and Gabeur, 2024) encapsulate transferable semantic priors and offer a promising path toward alleviating annotation dependency via transfer learning on downstream tasks like high-resolution RS image segmentation (Zhang et al., 2025).
-Although full fine-tuning of foundation models can enhance performance, it also incurs substantial computational and memory overheads (Han et al., 2024).
-Parameter-Efficient Fine-Tuning (PEFT) methods, such as Adapters (Houlsby and Giurgiu, 2019) and Low-Rank Adaptation (LoRA) (Hu and Shen, 2022), have emerged as viable alternatives for resource-conscious adaptation.
-However, due to the pronounced domain gap between the natural or medical images used for pre-training and the unique characteristics of RS data, conventional PEFT strategies often fail to generalize well in RS scenarios (Khanna and Irgau, 2024; Stojnic and Klemmer, 2025).
-Moreover, remote sensing imagery encompasses a wide spectrum of object scales—from large-scale vegetated regions to fine-grained human-made structures—necessitating robust multi-scale feature modeling capabilities (Cheng and Lei, 2023).
-To this end, we propose FreqWeaver, a frequency-temporal modulation adapter designed to enhance the adaptability of foundation models like SAM2 to ultra-high-resolution (UHR) remote sensing imagery.
-FreqWeaver decomposes input features into low- and high-frequency components via spectral filtering, which are then processed using large and small receptive field convolutions to respectively capture global semantics and fine-grained structures.
-A spatial-domain residual branch is further incorporated to complement missing local details.
-This design not only leverages the semantic priors encoded in SAM2 for basic object recognition, but also enhances the model’s capacity to represent diverse spatial patterns in UHR data.
-In particular, the proposed multi-scale feature interaction mechanism significantly improves the model’s ability to delineate fine boundary details of large-scale objects and accurately identify small, morphologically complex structures—thereby enabling precise and robust segmentation across heterogeneous geographic scenes under strict parameter constraints.
-Remote sensing (RS) systems, owing to their extensive spatial coverage and high temporal resolution, continuously collect vast amounts of unlabeled imagery, offering a rich yet underutilized resource for semi-supervised learning (SSL) (Yan and Yang, 2024).
-To address the challenges posed by limited supervision and high annotation costs in high-resolution RS segmentation, we propose a semi-supervised fine-tuning framework built upon a pre-trained foundation model.
-The framework integrates a teacher–student consistency learning paradigm with an uncertainty-aware weighting strategy, enabling efficient and robust domain adaptation.
-By dynamically modulating the contribution of pseudo-labels during training, the method effectively suppresses the adverse impact of noisy supervision—an issue particularly prevalent in RS imagery due to complex backgrounds, ambiguous boundaries, and high inter-class similarity.
-As a result, it significantly enhances model robustness and generalization across diverse geographic regions, acquisition conditions, and sensor types (Fan and Zhou, 2025), while simultaneously reducing reliance on costly manual annotations.
+Ultra-high Spatial Resolution (UHSR) remote sensing imagery, typically at sub-meter scales, offers rich spatial detail that allows accurate identification of individual ground objects and preservation of their structural features (Lv et al., 2023).
+Compared with coarser imagery, it provides sharper boundaries, finer textures, and the ability to capture small, heterogeneous features that are essential for detailed geospatial analysis (Guo et al., 2024).
+Leveraging these strengths, UHSR land cover classification converts high-detail imagery into thematic maps that combine broad spatial patterns with precise object-level information (Gui et al., 2024).
+The fine granularity of UHSR data reduces mixed-pixel effects, improves boundary accuracy, and enables reliable detection of subtle features such as narrow roads, isolated vegetation patches, and variations in roof materials or tree canopy structures.
+These advantages produce classification results that more closely match real-world surface conditions, supporting parcel-level land use mapping, high-precision urban modeling, ecosystem monitoring, and disaster impact assessment (Li et al., 2024a).
+By linking large-scale landscape observation with fine-scale object analysis, UHSR-based land cover classification has become an essential tool for modern geospatial applications.
+Despite its importance, most current remote sensing segmentation methods still focus on imagery with 1-meter spatial resolution, which is sufficient for identifying coarse land cover categories such as large vegetation patches, water bodies, and major transportation corridors, but struggles to capture fine-grained details such as narrow alleyways, small rooftop structures, or individual vehicles (Mendieta et al., 2024; Ma et al., 2025; Zhang et al., 2022).
+UHSR imagery, by contrast, enables the recognition of such detailed structures and subtle boundaries, offering the potential for parcel-level urban analysis and more precise land cover mapping (Huang et al., 2025).
+However, annotated UHSR datasets remain scarce and are extremely time- and labor-intensive to produce (Liu and Zhang, 2025; He et al., 2024), as pixel-level labeling at this resolution demands careful delineation of numerous small and heterogeneous objects.
+This annotation bottleneck has limited large-scale exploration of UHSR land cover classification in both research and practical applications.
+To address these limitations, we propose Baltimore Atlas, a novel framework for UHSR land cover classification.
+Baltimore Atlas performs multi-scale frequency-decoupled modeling for complex UHSR data patterns, preserving global boundaries while capturing fine-grained local details.
+Moreover, the framework leverages the prior knowledge of foundation models and integrates a semi-supervised learning strategy to reduce reliance on large annotated datasets.
+Consequently, the framework achieves high-accuracy and robust UHSR land cover mapping under limited training data availability.
+At the heart of Baltimore Atlas is the FreqWeaver Adapter, a frequency-modulated module designed to adapt SAM2 for UHSR imagery.
+To address the diverse spatial patterns in UHSR data, the module first applies spectral filtering to separate input features into low- and high-frequency components.
+The low-frequency branch employs large-receptive-field convolutions to capture global semantics, while the high-frequency branch applies small-receptive-field convolutions to preserve fine-grained structures.
+In parallel, a complementary spatial-domain residual branch restores local details.
+This design introduces only two additional convolutional kernels, maintaining a parameter-efficient architecture, and enables the model to effectively integrate global context with fine-grained detail for more accurate UHSR land cover classification.
+Further, we introduce the Uncertainty-Aware Teacher Student Framework, a semi-supervised framework for UHSR data.
+The teacher model takes weakly perturbed inputs and generates predictions.
+We compute pixel-wise entropy from these predictions, and lower entropy means higher confidence.
+For each pixel, the class with the highest confidence is chosen as the pseudo-label.
+High-confidence regions give stronger supervision to the student model, which is trained on strongly perturbed inputs.
+Focusing on these reliable regions helps the framework reduce noise from low-confidence areas, learn more stable representations, and improve performance and generalization in complex scenes.
+In addition, leveraging large-scale unlabeled data further reduces the reliance on annotated training samples.
 The main contributions of this paper are summarized as follows:
-  • PEFT-Driven Semi-Supervised Tuning of SAM2: We propose an Adapter-based PEFT strategy integrated with an uncertainty-guided consistency regularization mechanism for the SAM2 model (Ravi and Gabeur, 2024).
-This approach significantly reduces the dependency on labeled data, achieving efficient segmentation performance while updating fewer than $5.96\%$ of the total model parameters.
-  • 0.3m-Resolution Remote Sensing Dataset for Fine-Grained Segmentation: To support detailed land cover analysis, we construct a high-resolution (0.3m) remote sensing dataset over Baltimore.
-Compared to conventional 1m-Resolution datasets, it offers finer spatial granularity and greater object diversity, providing a more challenging and representative benchmark for evaluating segmentation models under real-world urban conditions.
-  • FreqWeaver Adapter for Disentangled Frequency–Temporal Multi-Scale Modeling: To better capture the fine-grained structures and scale-diverse patterns in high-resolution RS imagery, we introduce FreqWeaver, a frequency–temporal modulation adapter that decomposes features into low- and high-frequency components for separate processing via large and small receptive field convolutions.
-A complementary spatial-domain branch further enriches semantic representation.
-This design enhances SAM2’s capacity to model both global context and local detail, enabling more precise and parameter-efficient multi-scale segmentation.
-  • Uncertainty-Aware Semi-Supervised Learning: Considering the high error rates of pseudo-labels in remote sensing imagery—caused by complex backgrounds, fine-grained boundaries, and inter-class spectral similarity—we propose an uncertainty-aware semi-supervised framework that downweights unreliable predictions based on entropy.
-This design effectively suppresses misleading gradients from noisy supervision and improves the robustness of segmentation across diverse regions, sensors, and seasonal conditions.
+  • We propose Baltimore Atlas, a novel framework for UHSR land cover classification that achieves high-accuracy and robust mapping even with limited training data.
+  • We design a unified architecture with two core components: FreqWeaver Adapter, which parameter-efficiently adapts SAM2 to complex UHSR data patterns, and Uncertainty-Aware Teacher Student Framework, which uses entropy-based confidence to learn stable representations and reduce complex noise.
+  • We introduce the Baltimore Atlas Dataset, a 0.3 m resolution dataset based on aerial imagery of Baltimore City. Experiments on this dataset demonstrate consistent improvements over prior methods.
 2. Relate Work
-2.1. Foundation Models in Remote Sensing
-Foundation Models (FMs), notably Vision Transformer (ViT)-based architectures (Dosovitskiy et al., 2021), learn transferable representations from extensive unlabeled data, enabling strong generalization across diverse tasks.
-Adapting FMs to remote sensing (RS) is challenging due to RS imagery’s distinct characteristics (multi-scale spatial structures, rich spectral content, complex semantic distributions) that diverge significantly from natural images.
-To address these challenges, Geospatial Foundation Models (GeoFMs) have emerged.
-Models like Clay (Model, 2024) and Prithvi (Szwarcman et al., 2025), pre-trained on extensive satellite/aerial data, are optimized for Earth Observation (EO).
-Recent advances such as S2MAE (spatial-spectral pretraining) (Placeholder et al., 2024b), GeoChat (vision-language grounding) (Placeholder et al., 2024a), and SkySense (multi-modal integration) (Placeholder et al., 2024c) highlight the trend towards task-adaptive geospatial architectures.
-General-purpose FMs like the Segment Anything Model (SAM) (Kirillov et al., 2023b) and SAM2 (Ravi and Gabeur, 2024) offer prompt-driven, class-agnostic segmentation.
-However, the original SAM performs suboptimally on RS imagery, showing limited semantic awareness, difficulty with small/dense objects, and inconsistency in tasks like semantic change detection.
-These issues stem from its natural image bias and lack of RS domain priors.
-To address these SAM limitations, STSAMNet (Yang et al., 2025) integrates unsupervised domain adaptation (UDA), pseudo-label refinement, and adapter layers for improved building extraction.
-SCNet (Jiang et al., 2025) uses multi-scale feature fusion and semantic alignment to adapt SAM2 for semantic change detection.
-Prompt generation and CLIP-based supervision are also explored for semantic guidance in RS.
-These adaptations show a shift from universal zero-shot inference to domain-specialized fine-tuning.
-Thus, RS is not just adopting FMs but also shaping their development.
-This is driven by RS challenges like scale variation, semantic ambiguity, and the need for label-efficient geospatial analysis.
-2.2. Semantic Segmentation for Complex Remote Sensing
-Remote sensing image segmentation faces unique challenges.
-Pixel-level annotation scarcity is primary, severely limiting fully supervised methods.
-Second, inherent multi-scale spatial heterogeneity necessitates models adept at handling drastic scale variations (small objects to vast regions).
-Furthermore, high inter-class similarity and significant intra-class diversity (from varying illumination, seasons, and sensors) severely test model feature discriminability.
-Semi-Supervised Semantic Segmentation (SSSS) addresses annotation scarcity by leveraging abundant unlabeled data.
-Key SSSS strategies include consistency regularization and pseudo-labeling(Tarvainen and Valpola, 2018; Chen et al., 2021; Sohn and Berthelot, 2020).
-For instance, MUCA (Wang et al., 2025) uses a Multi-scale Uncertainty Consistency (MSUC) module and cross-teacher-student attention to address multi-scale objects and inter-class similarity, enhancing multi-layer semantic consistency and discriminability.
-SegKC (Than et al., 2025) improves pseudo-label quality and efficiency via heterogeneous dual-network co-training with bi-directional knowledge transfer, using a lightweight network for efficient inference.
-Unsupervised Domain Adaptation (UDA) addresses domain shifts in remote sensing data, often from varying sensors or locations.
-For example, STSAMNet (Yang et al., 2025) combines UDA with the Segment Anything Model (SAM) via an adapter, optimizing segmentation for target-domain high-resolution imagery and mitigating performance drops in new regions.
-While SSSS and UDA methods partially address label scarcity and domain shifts, accurately segmenting complex spatial structures and multi-scale targets remains a bottleneck.
-Many methods struggle to capture diverse multi-scale features or introduce artifacts during pseudo-labeling and domain adaptation, especially at boundaries or in textured areas.
-To overcome these limitations, we explore parameter-efficient solutions that not only reduce reliance on labeled data but also enhance multi-scale representation under minimal supervision, as detailed next.
-2.3. Parameter-Efficient Fine-Tuning Method
-As foundation models (FMs), including geospatial variants (GeoFMs), continue to grow in size and complexity, full fine-tuning has become increasingly impractical due to high computational costs and risks such as overfitting and catastrophic forgetting.
-Parameter-Efficient Fine-Tuning (PEFT) offers a scalable alternative by adapting models through the selective update of a small subset of parameters, while keeping the majority of pre-trained weights frozen.
-Several PEFT techniques have been proposed to balance efficiency and performance.
-Adapter tuning, which inserts lightweight modules into transformer layers, has been successfully applied in models like STSAMNet (Yang et al., 2025) for adapting SAM to high-resolution remote sensing.
-Low-Rank Adaptation (LoRA) (Hu et al., 2021) reduces parameter overhead by constraining updates to low-rank subspaces.
-Prompt tuning (Jia et al., 2022) and prefix tuning (Li and Liang, 2021) modify task-specific behavior through learned embeddings, without altering the base model.
-These methods significantly reduce training cost and memory usage, and often enhance generalization by preserving pre-trained representations.
-PEFT has gained traction in remote sensing.
-(Marti-Escofet et al., 2025) benchmarked multiple PEFT methods across diverse GeoFM architectures and Earth observation datasets, showing that PEFT can match or surpass full fine-tuning with far lower resource requirements.
-(He et al., 2023) introduced Sensitivity-aware PEFT (SPT), which identifies and prioritizes task-relevant parameters for adaptation, integrating LoRA or adapters accordingly.
-Domain-specific approaches such as AiRs (Hu et al., 2024b), TEA (Hu et al., 2024a), and Earth-Adapter (Hu et al., 2025) further enhance transformer adaptation for Earth observation via residual-enhanced adapters.
-To address key challenges in remote sensing imagery—such as large parameter overhead, aliasing artifacts, and significant object-scale variability—we propose a novel adapter architecture that combines frequency-domain decomposition with specialized convolutional processing.
-By disentangling high- and low-frequency components through dedicated convolutional branches and incorporating temporal information, the proposed method enables effective multi-scale adaptation while maintaining minimal parameter cost.
-Figure 1. Overview of the proposed framework.
+2.1. Ultra-high Spatial Resolution Land Cover Classification
+Early land cover classification mainly used pixel-based methods, treating each pixel as an independent unit classified only by its spectral features.
+In high spatial resolution imagery, these methods often cause confusion and fragmentation due to the high spectral similarity among different land covers (Roy et al., 2018; Han et al., 2020), leading to the development of Geographic Object-Based Image Analysis (GEOBIA/OBIA) (Chen et al., 2018; Hay and Castilla, 2008; Kucharczyk et al., 2020).
+GEOBIA aggregates similar pixels into image objects and combines segmentation, feature extraction, and semantic rules to improve classification, but large-scale mapping still requires considerable expert knowledge and manual intervention.
+To reduce manual effort and improve automation, deep learning has been widely applied to fine-grained land cover classification.
+LSHR-LCM (Robinson et al., 2019) evaluated multiple deep learning models on 1 m resolution data across the United States, demonstrating the feasibility of convolutional networks for nationwide mapping but showing strong dependence on high-quality labeled data.
+ABCNet (Li et al., 2021) improved network design to preserve fine spatial details while capturing global context, enhancing both accuracy and stability in high-resolution semantic segmentation.
+Urban Watch (Zhang et al., 2022) refined the impervious surface class, resulting in a total of nine urban categories, to address material diversity and overly coarse definitions, and incorporated semi-supervised learning to reduce annotation requirements.
+LEG (Li et al., 2024b) used 30 m low-resolution labels for weak supervision, improving 1 m mapping accuracy and demonstrating the potential for cross-resolution transfer.
+TLULC (Khan et al., 2024) applied large-scale Transformer pretraining and transfer learning, further enhancing the generalization and robustness of high-resolution mapping.
+Despite these advances, most methods focus on 1 m or coarser resolutions, leaving UHSR land cover classification largely unexplored.
+To address this gap, we propose the Baltimore Atlas, a framework for UHSR(0.3 m resolution) land cover classification that combines tailored transfer learning with semi-supervised learning to adapt to the complex spatial patterns of UHSR imagery, achieving improved performance while reducing reliance on large-scale annotated datasets.
+2.2. Parameter-Efficient Adaptation for Multi Scale Remote Sensing Understanding
+Foundation models (FMs) such as SAM2 (Ravi et al., 2024) are trained on large-scale natural image datasets with segmentation labels, enabling strong generalization across segmentation tasks.
+DED-SAM (Qiu et al., 2025) demonstrates SAM2’s zero-shot segmentation capability on remote sensing (RS) imagery.
+To better align model representations with RS characteristics, GeoFMs such as Clay (Model, 2024), Prithvi (Szwarcman et al., 2024), and S2MAE (Li et al., 2024c) incorporate RS-specific priors through pretraining on satellite and aerial imagery.
+However, these models still lack task-specific priors for segmentation.
+As a result, both general-purpose FMs and GeoFMs require adaptation to downstream datasets to achieve strong performance in tasks like land cover classification.
+Parameter-efficient fine-tuning (PEFT) addresses this need by reducing trainable parameters while maintaining accuracy: LoRA (Hu et al., 2022) constrains weight updates to low-rank subspaces, prompt tuning (Jia et al., 2022) and prefix tuning (Li and Liang, 2021) insert learnable vectors into the input or attention layers, and adapter tuning (Chen et al., 2022) adds lightweight modules inside transformer layers.
+While effective in many domains, these approaches often struggle with the multi-scale structures, rich spectral content, and complex semantics of RS data.
+To bridge this gap, RS-specific PEFT designs have been proposed: AiRs (Hu et al., 2024b) models spatial context and semantic responses, TEA (Hu et al., 2024a) employs a lightweight side adapter with macro-scale semantic guidance, and Earth-Adapter (Hu et al., 2025) applies frequency decoupling for multi-scale modeling but relies on a unified linear encoder, limiting scale-specific feature modeling.
+To address challenges in remote sensing imagery, we propose the FreqWeaver Adapter, a novel adapter design that integrates frequency-domain decomposition with convolutional processing.
+It separates high- and low-frequency components through dedicated convolutional branches and incorporates spatial information.
+This design enables efficient multi-scale adaptation, reduces aliasing artifacts, avoids large variations in object scale, while maintaining a very low additional parameter cost.
+2.3. Semi-Supervised Learning for Robust Remote Sensing Representations
+Semi-supervised learning (SSL) leverages both labeled and unlabeled data to improve model generalization and robustness under limited annotations.
+General SSL strategies include pseudo-labeling (Lee et al., 2013; Xie et al., 2020), consistency regularization (Berthelot et al., 2019; Sohn and Berthelot, 2020), and teacher student frameworks (Tarvainen and Valpola, 2017).
+While these approaches have demonstrated strong performance in natural image tasks, their direct application to remote sensing (RS) is often limited by challenges such as multi-scale spatial structures and high inter-class spectral similarity.
+To overcome these limitations, several methods have extended SSL to better fit RS characteristics: SSSM (Feng et al., 2025) introduced multi-level feature fusion with pseudo-label sampling for scene classification, alleviating the noise caused by low-confidence labels; UniMatch (Yang et al., 2023) enhanced semantic segmentation with task-driven perturbation branches to better handle complex RS structures; MUCA (Wang et al., 2025) incorporated multi-scale uncertainty consistency and cross teacher–student attention to align intermediate features across scales, improving the discrimination of spectrally similar land-cover classes; and SOOD (Hua et al., 2023) adapted semi-supervised learning to oriented object detection by introducing orientation-consistency and global-consistency losses, significantly improving detection performance on aerial benchmarks.
+Overall, these methods highlight the potential of SSL to reduce annotation costs and enhance performance in complex RS scenarios, but their applications to UHSR data remain insufficiently explored.
+To reduce the noise caused by unstable predictions in UHSR imagery, we propose the Uncertainty-Aware Teacher–Student Framework.
+This framework uses entropy-based uncertainty estimation to select high-confidence pseudo-labels and focuses supervision on reliable regions.
+In this way, it limits the effect of noisy predictions, makes better use of large amounts of unlabeled data, and learns more stable feature representations while lowering dependence on extensive annotations.
+Figure 1.
+Overview of the proposed framework.
+Given a remote sensing image, two augmented views are generated: a strongly augmented version is fed into the student network, and a weakly augmented version is processed by the teacher network.
+To capture the characteristics of UHSR data, a parameter-efficient FreqWeaver Adapter is inserted into specific Transformer layers.
+The teacher network is updated with an exponential moving average (EMA) strategy to maintain stable parameter evolution.
+For unlabeled data, the teacher network provides pseudo-labels to enforce consistency with student predictions, enabling more robust feature learning.
+For labeled data, ground-truth annotations directly supervise the student network to ensure accurate semantic understanding.
+Overall, the framework integrates adapter-based adaptation with semi-supervised learning to deliver reliable segmentation of UHSR imagery while reducing reliance on labeled data.
 3. Method
-We propose a semi-supervised, parameter-efficient fine-tuning framework built upon the Segment Anything Model 2 (SAM2), which effectively leverages the prior knowledge of large-scale foundation models.
-To this end, we introduce the FreqWeaver Adapter—a lightweight yet effective module that adapts SAM2 to remote sensing data while maintaining high parameter efficiency.
-In addition, we design a semi-supervised learning strategy that fully utilizes unlabeled data to improve domain adaptation and generalization under limited supervision.
-The overall framework is illustrated in Figure 1.
-3.1. FreqWeaver: Frequency-Modulated Adapter for Efficient RS Fine-Tuning
-Remote sensing (RS) imagery is typically characterized by significant scale variation, densely distributed objects, and complex textures.
-These properties pose challenges not only to general-purpose vision models such as SAM, but also to existing parameter-efficient tuning methods, which often struggle to adapt to such complex data distributions.
-To address this, we propose FreqWeaver, a lightweight frequency–temporal modulation adapter.
-FreqWeaver explicitly decomposes input features into low- and high-frequency components, which are processed separately using convolutional filters with distinct receptive fields to model global semantics and fine-grained details, respectively.
-In addition, a temporal branch is introduced to capture residual semantic information that may not be effectively represented in the frequency domain.
-This design significantly improves the adaptability and structural representation capability of SAM on RS imagery, while maintaining parameter efficiency.An overview of the proposed adapter is illustrated in Figure 2.
-3.1.1. Spectral Decomposition
-To facilitate frequency-aware feature refinement, we perform spectral decomposition on the encoder outputs using the 2D Discrete Fourier Transform (DFT).
-This allows explicit separation of low- and high-frequency information for targeted processing.
-Before applying the DFT, we first reshape the input feature tensor from $\mathbf{X}\in\mathbb{R}^{B\times H\times W\times C}$ to $\mathbf{X}^{\prime}\in\mathbb{R}^{B\times C\times H\times W}$ to align with the channel-wise Fourier transform convention.
+3.1. Overview
+We propose Baltimore Atlas, a framework for ultra-high spatial resolution (UHSR) land cover classification.
+To support this framework, we construct the Baltimore Atlas Dataset, a benchmark created from 0.3 m aerial imagery of Baltimore City acquired through the National Agriculture Imagery Program (NAIP).
+For the model, we adopt SAM2 as the foundation and introduce the FreqWeaver Adapter, which enables parameter-efficient fine-tuning and adapts SAM2 to the multi-scale patterns of UHSR data.
+In addition, we design the Uncertainty-Aware Teacher–Student Framework, a semi-supervised learning method that employs uncertainty estimation to generate high-confidence pseudo-labels, which are then used as supervisory signals for training on unlabeled data.
+Overall, the integration of these components enables accurate land cover classification on UHSR data with limited annotated samples, as illustrated in Fig.
+1.
+3.2. FreqWeaver Adapter
+A key challenge of UHSR imagery lies in its significant scale variation, dense object distribution, and complex textures, which limit the adaptability of SAM2 and existing parameter-efficient tuning (PEFT) methods.
+To address this, we introduce FreqWeaver, a lightweight frequency–spatial modulation adapter that improves multi-scale modeling while keeping parameter cost low.
+The method integrates frequency-domain decomposition with convolutional processing: it separates low- and high-frequency components through dedicated convolutional branches and incorporates spatial information for more complete feature representation.
+Our approach is organized into three stages—Frequency Decomposition, Multi-Branch Feature Modeling, and Adaptive Feature Fusion—which will be described in the following sections.
+An overview of the proposed adapter is illustrated in Figure 2.
+3.2.1. Frequency Decomposition
+To enable frequency-aware feature refinement, we apply frequency decomposition to the encoder features using the 2D Discrete Fourier Transform (DFT).
+This operation explicitly separates low- and high-frequency components for specialized processing.
+Before applying the DFT, the input feature tensor is reshaped from $\mathbf{X}\in\mathbb{R}^{B\times H\times W\times C}$ to $\mathbf{X}^{\prime}\in\mathbb{R}^{B\times C\times H\times W}$ to align with the channel-wise Fourier transform convention.
 We then apply the 2D DFT independently to each channel as:
@@
 The resulting spectrum $\mathbf{X}_{\mathrm{f}}\in\mathbb{C}^{B\times C\times H\times W}$ contains complex-valued frequency coefficients.
-To extract different frequency bands, we define two complementary radial binary masks:
+To extract different frequency bands, two complementary radial binary masks are defined:
 (2) $\mathcal{M}_{\mathrm{lf}}=\mathbb{I}(r<\rho),\quad\mathcal{M}_{\mathrm{hf}}=1-\mathcal{M}_{\mathrm{lf}},$
-where $r$ denotes the radial distance from the spectral center, and $\rho$ is a predefined cutoff threshold.
-We isolate the low- and high-frequency components by applying the masks element-wise:
+where $\mathbb{I}(\cdot)$ denotes the indicator function, $r$ is the radial distance from the spectral center, and $\rho$ is a predefined cutoff threshold.
+The low- and high-frequency components are isolated by applying the masks element-wise:
 (3) $\mathbf{X}_{\mathrm{lf}}=\mathbf{X}_{\mathrm{f}}\odot\mathcal{M}_{\mathrm{lf}},\quad\mathbf{X}_{\mathrm{hf}}=\mathbf{X}_{\mathrm{f}}\odot\mathcal{M}_{\mathrm{hf}},$
 where $\odot$ denotes Hadamard (element-wise) multiplication.
-Finally, we recover the spatial-domain representations via the inverse Fourier transform:
+Finally, the spatial-domain representations are recovered via the inverse Fourier transform:
 (4) $\widehat{\mathbf{X}}_{\mathrm{lf}}=\mathcal{F}^{-1}_{\mathrm{DFT}}(\mathbf{X}_{\mathrm{lf}}),\quad\widehat{\mathbf{X}}_{\mathrm{hf}}=\mathcal{F}^{-1}_{\mathrm{DFT}}(\mathbf{X}_{\mathrm{hf}}),$
-resulting in real-valued tensors $\widehat{\mathbf{X}}_{\mathrm{lf}},\widehat{\mathbf{X}}_{\mathrm{hf}}\in\mathbb{R}^{B\times C\times H\times W}$ representing the low- and high-frequency spatial features, respectively.
-These components are subsequently passed to dedicated branches for frequency-specific refinement.
-Figure 2. Overview of the proposed FreqWeaver Adapter.
-3.1.2. Frequency–Temporal Feature Modulation
-To enhance frequency- and spatial-domain representations in a disentangled manner, FreqWeaver introduces three specialized branches, each responsible for modeling a specific component of the feature spectrum: low-frequency (global structure), high-frequency (local detail), and spatial-domain semantics.
+yielding real-valued tensors $\widehat{\mathbf{X}}_{\mathrm{lf}},\widehat{\mathbf{X}}_{\mathrm{hf}}\in\mathbb{R}^{B\times C\times H\times W}$ that represent the low- and high-frequency spatial features.
+These components are then forwarded to separate branches for frequency-specific refinement.
+Figure 2.
+Overview of the proposed FreqWeaver Adapter.
+The FreqWeaver Adapter is inserted into specific Transformer layers as the only trainable component, while all other parameters remain frozen.
+It performs frequency decoupling by applying a Fourier transform to separate high- and low-frequency components.
+High-frequency information is processed with small convolution kernels to capture local textures and fine details, while low-frequency information is modeled with large kernels to enhance global boundary features.
+To prevent information loss, the original spatial features are preserved, and a router is employed to fuse multi-scale information effectively.
+This design provides more comprehensive feature representations and enables the model to better adapt to the multi-scale characteristics of UHSR data.
+3.2.2. Multi-Branch Feature Modeling
+To enhance frequency- and spatial-domain representations in a disentangled manner, FreqWeaver introduces three specialized branches, each responsible for modeling a specific component of the feature representations: low-frequency (global structure), high-frequency (local detail), and spatial-domain semantics.
 These branches operate independently and produce complementary signals for subsequent fusion.
@@
 The low-frequency response is computed as:
-(5) $\mathbf{F}_{\mathrm{lf}}=\mathcal{F}_{\mathrm{LF}}(\widehat{\mathbf{X}}_{\mathrm{lf}}),$
-where $\widehat{\mathbf{X}}_{\mathrm{lf}}\in\mathbb{R}^{B\times C\times H\times W}$ is the input low-frequency feature map, and $\mathcal{F}_{\mathrm{LF}}(\cdot)$ denotes a depthwise convolution with a kernel size of $11\times 1$ 11, initialized with zeros.
-The output $\mathbf{F}_{\mathrm{lf}}\in\mathbb{R}^{B\times C\times H\times W}$ encodes coarse structural information and is reshaped to $\mathbb{R}^{B\times H\times W\times C}$.
+(5) $\mathbf{F}_{\mathrm{lf}}=f_{\mathrm{lf}}(\widehat{\mathbf{X}}_{\mathrm{lf}}),$
+where $\widehat{\mathbf{X}}_{\mathrm{lf}}\in\mathbb{R}^{B\times C\times H\times W}$ is the input low-frequency feature map, and $f_{\mathrm{lf}}(\cdot)$ denotes a depthwise convolution with a kernel size of $11\times 11$.
+The output $\mathbf{F}_{\mathrm{lf}}\in\mathbb{R}^{B\times C\times H\times W}$ encodes coarse structural information and is reshaped into $\mathbb{R}^{B\times H\times W\times C}$.
 High-Frequency Branch.
@@
 It enhances boundary clarity and preserves local precision.
-The high-frequency response is defined as:
-(6) $\mathbf{F}_{\mathrm{hf}}=\mathcal{F}_{\mathrm{HF}}(\widehat{\mathbf{X}}_{\mathrm{hf}}),$
-where $\widehat{\mathbf{X}}_{\mathrm{hf}}\in\mathbb{R}^{B\times C\times H\times W}$ is the high-frequency input, and $\mathcal{F}_{\mathrm{HF}}(\cdot)$ is a depthwise convolution with a $5\times 5$ kernel.
-The output $\mathbf{F}_{\mathrm{hf}}\in\mathbb{R}^{B\times C\times H\times W}$ highlights spatially localized structures and is similarly reshaped to $\mathbb{R}^{B\times H\times W\times C}$.
+The high-frequency response is computed as:
+(6) $\mathbf{F}_{\mathrm{hf}}=f_{\mathrm{hf}}(\widehat{\mathbf{X}}_{\mathrm{hf}}),$
+where $\widehat{\mathbf{X}}_{\mathrm{hf}}\in\mathbb{R}^{B\times C\times H\times W}$ is the high-frequency input, and $f_{\mathrm{hf}}(\cdot)$ is a depthwise convolution with a $5\times 5$ kernel.
+The output $\mathbf{F}_{\mathrm{hf}}\in\mathbb{R}^{B\times C\times H\times W}$ highlights spatially localized structures and is reshaped into $\mathbb{R}^{B\times H\times W\times C}$.
 Spatial Branch.
-This branch directly processes the spatial-domain feature without spectral filtering, enabling the modeling of mid-frequency signals and ambiguous semantic cues that may not be cleanly separated in the frequency domain.
-It plays a complementary role in capturing transitional or occluded content.
-The spatial-domain response is formulated as:
-(7) $\mathbf{F}_{\mathrm{s}}=\mathcal{F}_{\mathrm{S}}(\mathbf{X}),$
-where $\mathbf{X}\in\mathbb{R}^{B\times H\times W\times C}$ is the original feature map before frequency decomposition.
-The operator $\mathcal{F}_{\mathrm{S}}(\cdot)$ is a two-layer bottleneck MLP.
-The output $\mathbf{F}_{\mathrm{s}}\in\mathbb{R}^{B\times H\times W\times C}$ captures spatial-domain semantics that complement the frequency-based branches.
-3.1.3. Content-Aware Feature Fusion
-To integrate the outputs of the frequency–temporal branches, we adopt a content-aware fusion strategy guided by global contextual information.
-This mechanism allows the model to dynamically balance the relative contributions of low-frequency, high-frequency, and spatial-domain features on a per-sample basis.
-We first compute a global context descriptor via global average pooling over the spatial dimensions:
+This branch directly processes the spatial-domain features without frequency filtering.
+It is designed to recover residual semantics that may be overlooked by the low- and high-frequency branches, such as transitional regions, occluded objects, or mixed textures.
+The spatial-domain response is defined as:
+(7) $\mathbf{F}_{\mathrm{s}}=f_{\mathrm{s}}(\mathbf{X}),$
+where $\mathbf{X}\in\mathbb{R}^{B\times H\times W\times C}$ is the original feature map before frequency decomposition, and $f_{\mathrm{s}}(\cdot)$ is a two-layer bottleneck MLP.
+The output $\mathbf{F}_{\mathrm{s}}\in\mathbb{R}^{B\times H\times W\times C}$ provides complementary spatial semantics to the frequency-based branches.
+3.2.3. Adaptive Feature Fusion
+To integrate the outputs of the low-frequency, high-frequency, and spatial branches, we adopt a content-aware fusion strategy guided by global contextual information.
+This mechanism enables the model to dynamically balance the relative contributions of the three feature types on a per-sample basis.
+We first compute a global context descriptor by performing global average pooling over the spatial dimensions:
 (8) $\mathbf{g}=\frac{1}{HW}\sum_{h=1}^{H}\sum_{w=1}^{W}\mathbf{X}_{:,h,w,:}\in\mathbb{R}^{B\times C},$
-where $\mathbf{X}\in\mathbb{R}^{B\times H\times W\times C}$ is the original spatial-domain feature map.
+where $\mathbf{X}\in\mathbb{R}^{B\times H\times W\times C}$ denotes the original spatial-domain feature map.
 The descriptor $\mathbf{g}$ is then passed to a lightweight router MLP, which outputs normalized fusion weights:
-(9) $[r_{\mathrm{s}},r_{\mathrm{lf}},r_{\mathrm{hf}}]=\operatorname{softmax}(\mathcal{F}_{\mathrm{Router}}(\mathbf{g}))\in\mathbb{R}^{B\times 3},$
-where $\mathcal{F}_{\mathrm{Router}}(\cdot)$ denotes a two-layer MLP with softmax normalization to ensure $\sum_{k}r_{k}=1$.
-Finally, the three modulation outputs are adaptively fused with the original features via weighted addition:
+(9) $r_{\mathrm{s}},r_{\mathrm{lf}},r_{\mathrm{hf}}]=\operatorname{softmax}(f_{\mathrm{router}}(\mathbf{g}))\in\mathbb{R}^{B\times 3},$
+where $f_{\mathrm{router}}(\cdot)$ denotes a two-layer MLP with softmax normalization to ensure $\sum_{k}r_{k}=1$.
+Finally, the three branch outputs are fused with the original features via weighted addition:
 (10) $\mathbf{Y}=\mathbf{X}+r_{\mathrm{s}}\odot\mathbf{f}_{\mathrm{s}}+r_{\mathrm{lf}}\odot\mathbf{f}_{\mathrm{lf}}+r_{\mathrm{hf}}\odot\mathbf{f}_{\mathrm{hf}},$
@@
 The final output $\mathbf{Y}\in\mathbb{R}^{B\times H\times W\times C}$ preserves the original spatial resolution while incorporating frequency- and content-aware enhancements.
-This fusion strategy enables adaptive emphasis of complementary information sources, thereby improving the model’s ability to generalize to diverse spatial patterns in remote sensing imagery.
-3.2. Uncertainty-Aware Teacher–Student Adaptation
-In the context of remote sensing segmentation, acquiring large-scale annotated datasets is both expensive and labor-intensive.
-To alleviate this constraint, we introduce an Uncertainty-Aware Mean Teacher (UAMT) framework that leverages unlabeled data via a dual-network consistency paradigm.
-The method enhances pseudo-supervision reliability by integrating entropy-based uncertainty estimation into a perturbation-consistent learning scheme.
-This design allows for stable adaptation of large vision foundation models (e.g., SAM2) with minimal supervision.
+This fusion strategy adaptively balances complementary information sources, thereby improving the model’s ability to generalize to diverse spatial patterns in remote sensing imagery.
+3.3. Uncertainty-Aware Teacher–Student Framework
+To address the noise in pseudo-labels for semi-supervised remote sensing segmentation, especially in UHSR imagery, we propose the Uncertainty-Aware Mean Teacher (UAMT) framework.
+The method employs entropy-based uncertainty estimation to select high-confidence pseudo-labels as supervisory signals for unlabeled data, reducing the impact of uncertain predictions.
+By combining this mechanism with a perturbation-consistent teacher–student scheme, UAMT stabilizes pseudo-supervision and enables efficient adaptation of SAM2 under limited annotation.
 3.3.1. Mean Teacher Framework with Perturbation Consistency
@@
 3.3.2. Entropy-Based Uncertainty Estimation
-While teacher predictions provide valuable supervisory cues, pseudo-labels generated for remote sensing images are particularly error-prone due to complex backgrounds, fine-grained object boundaries, and domain-specific variations.
-To mitigate the impact of such noisy supervision, we employ a pixel-wise confidence estimation mechanism based on Shannon entropy to downweight uncertain predictions.
-Let $\mathbf{P}^{\text{T}}_{j}\in\mathbb{R}^{C}$ denote the softmax probability vector predicted by the teacher at pixel $j$, where $C$ is the number of classes.
-The uncertainty score is quantified by entropy:
+Pseudo-labels in remote sensing segmentation are often noisy due to complex backgrounds, fine boundaries, and domain-specific variations.
+To reduce their negative impact, we apply pixel-wise confidence weighting based on Shannon entropy.
+Let $\mathbf{P}^{\text{T}}_{j}\in\mathbb{R}^{C}$ be the teacher’s softmax probability vector at pixel $j$, where $C$ is the number of classes. The uncertainty is measured by entropy:
 (12) $\mathcal{H}_{j}=-\sum_{c=1}^{C}\mathbf{P}^{\text{T}}_{j,c}\log\mathbf{P}^{\text{T}}_{j,c},$
-which attains its maximum at $\log C$ when the prediction is maximally ambiguous (i.e., uniform distribution). To normalize this measure and convert it into a usable confidence weight, we define:
+which reaches $\log C$ when the distribution is uniform. We normalize this score and convert it to a confidence weight:
 (13) $w_{j}=1-\frac{\mathcal{H}_{j}}{\log C},\quad w_{j}\in[0,1],$
-where higher $w_{j}$ values indicate more confident predictions and hence stronger training signals.
-This design enables the model to suppress the impact of high-entropy (uncertain) regions and focus on reliable pseudo-labels.
+where higher $w_{j}$ values indicate more reliable predictions and thus stronger supervision. This scheme downweights uncertain regions and emphasizes high-confidence pseudo-labels.
 3.3.3. Loss Function: Supervised–Unsupervised Coupling
-The overall objective consists of two complementary terms: a supervised loss on labeled data and a confidence-weighted consistency loss on unlabeled data.
+The training objective combines a supervised loss on labeled data and a confidence-weighted consistency loss on unlabeled data.
 Supervised Loss.
-Given labeled samples $(x_{l},y_{l})\in\mathcal{D}_{l}$, the student model is trained with standard pixel-wise cross-entropy:
+For labeled pairs $(x_{l},y_{l})\in\mathcal{D}_{l}$, the student is trained with pixel-wise cross-entropy:
 (14) $\mathcal{L}_{\mathrm{sup}}=\mathrm{CE}\!\left(f(x_{l};\boldsymbol{\theta}_{\mathrm{S}}),y_{l}\right),$
 where $f(\cdot;\boldsymbol{\theta}_{\mathrm{S}})$ denotes the student prediction.
-Consistency Loss with Confidence Weighting.
-For unlabeled images $x_{u}\in\mathcal{D}_{u}$, the teacher generates soft predictions $\mathbf{P}^{\text{T}}$, while the student predicts $\mathbf{P}^{\text{S}}$ from a strongly perturbed view.
-A pseudo-label $\hat{y}^{\text{T}}_{j}=\arg\max_{c}\mathbf{P}^{\text{T}}_{j,c}$ is assigned per pixel, and the confidence-weighted consistency loss is formulated as:
-(15) $\mathcal{L}_{\mathrm{cons}}=\frac{1}{|\mathcal{D}_{u}|\cdot N}\sum_{x_{u}\in\mathcal{D}_{u}}\sum_{j=1}^{N}w_{j}\cdot\mathrm{CE}(\mathbf{P}^{\text{S}}_{j},\hat{y}^{\text{T}}_{j}),$
-where $N$ is the total number of pixels per image, and $\mathbf{P}^{\text{S}}_{j}$ is the predicted softmax from the student at pixel $j$.
+Consistency Loss.
+For unlabeled inputs $x_{u}\in\mathcal{D}_{u}$, the teacher produces probabilities $\mathbf{P}^{\text{T}}$, from which a pseudo-label $\hat{y}^{\text{T}}_{j}=\arg\max_{c}\mathbf{P}^{\text{T}}_{j,c}$ is obtained per pixel.
+The student predicts $\mathbf{P}^{\text{S}}$ from a perturbed view.
+Using confidence weights $w_{j}$ derived from entropy, the consistency loss is:
+(15) $\mathcal{L}_{\mathrm{cons}}=\frac{1}{B\cdot N}\sum_{x_{u}\in\mathcal{D}_{u}}\sum_{j=1}^{N}w_{j}\cdot\mathrm{CE}(\mathbf{P}^{\text{S}}_{j},\hat{y}^{\text{T}}_{j}),$
+where $B$ is the batch size and $N$ is the number of pixels per image.
 Total Objective.
-The combined loss is given by:
-(16) $\mathcal{L}_{\mathrm{total}}=\mathcal{L}_{\mathrm{sup}}+\lambda\cdot\mathcal{L}_{\mathrm{cons}},$
-where $\lambda=0.1$ balances the contributions of labeled and unlabeled samples.
-During training, only the parameters of the FreqWeaver adapters are updated, while the SAM2 backbone remains frozen.
-This strategy ensures efficient and stable transfer to remote sensing domains, leveraging both labeled and unlabeled data in a principled manner.
-Figure 3. Visualization of the State-of-the-Art PEFT Method.
+The final loss is:
+(16) $\mathcal{L}_{\mathrm{total}}=\mathcal{L}_{\mathrm{sup}}+\lambda\mathcal{L}_{\mathrm{cons}},$
+with $\lambda=0.1$ controlling the balance between labeled and unlabeled data.
+During training, only the parameters of FreqWeaver adapters are updated, while the SAM2 backbone remains frozen, ensuring efficient transfer to remote sensing domains.
 4. Experiment
 4.1. Dataset
-In this study, all experiments were conducted on ultra-high-resolution remote sensing imagery of Baltimore, captured at a spatial resolution of 0.3 meters.
-A total of 60 images, each with dimensions of 1024×1024 pixels, were selected based on the nine urban semantic categories defined by UrbanWatch, ensuring a relatively balanced distribution of class labels, as shown in Fig 4.
-The dataset comprises a total of 62,914,560 pixels.For initial annotations, we utilized the UrbanWatch model, which was originally trained on 1-meter resolution imagery.
-Although not optimized for ultra-high-resolution inputs, this model offered a coarse yet informative baseline.
-To improve annotation accuracy, we subsequently performed meticulous manual corrections.
-Given the complexity of remote sensing scenes and the difficulty of pixel-level labeling, this process was both labor-intensive and time-consuming.For training and evaluation, we used 55 images for training and the remaining 5 for testing.
-The test set alone contained 5,242,880 pixels, offering a sufficiently large sample size for robust and reliable performance evaluation.
+In this study, we introduce the Baltimore Atlas Dataset, an ultra-high-resolution remote sensing segmentation dataset constructed from 0.3 m aerial imagery of Baltimore City acquired through the National Agriculture Imagery Program (NAIP).
+The imagery is annotated into nine urban semantic categories defined by UrbanWatch through a two-stage process.
+In the first stage, we use the UrbanWatch model, trained on 1 m resolution imagery, to generate the initial annotations as a coarse baseline.
+In the second stage, these annotations are refined through manual corrections to improve accuracy.
+For sampling, we extract 1024 × 1024 pixel patches (300 m × 300 m) from the entire imagery to meet SAM2’s input requirements, while applying uniform sampling to maintain class balance.
+This process yields 60 samples (62,914,560 pixels in total), with the class distribution shown in Fig.
+3.
+Among them, 55 are used for training and 5 for testing, providing enough scale and diversity for effective model training and reliable performance evaluation.
 Figure 3. Visualization of land cover class distribution.
+Figure 4.
+Visualization of the State-of-the-Art PEFT Method.
+In the first row, our method shows better capture of textures and boundaries of small objects in dense urban areas, such as residential swimming pools.
+In the second and third rows, parking lots are recognized at the level of individual vehicles.
+Although this differs from region-level annotations, it highlights the strength of our local detail modeling and suggests directions for future fine-grained classification.
 4.2. Evaluation Metrics
 To comprehensively evaluate the segmentation performance, we adopt two widely used metrics: Intersection over Union (IoU) and Dice coefficient.
-Both metrics quantify the region-level overlap between predicted segmentation results and ground truth annotations, providing a general assessment of segmentation accuracy.
+Both metrics quantify the region-level overlap between predicted segmentation results and ground truth annotations, but with different emphases: IoU focuses on overall overlap consistency, while Dice is more sensitive to small objects and class imbalance.
+In combination, these metrics provide a comprehensive and complementary assessment of segmentation accuracy.
 Given the predicted region $P$ and the ground truth region $G$, the IoU and Dice coefficient are defined as follows:
@@
 4.3. Training and Testing Details
-All experiments were conducted using PyTorch on a single NVIDIA RTX L40S GPU with 48 GB of memory.
+All experiments are conducted using PyTorch on a single NVIDIA RTX L40S GPU with 48 GB of memory.
 In our work, we adopt SAM2 as the backbone network.
-Input images were uniformly resized to $1024\times 1024$ during both training and testing.
-The model was trained for 200 epochs using the Adam optimizer with a learning rate of $5\times 10^{-4}$ and a batch size of 2.
-For the segmentation task, we employed the standard Cross-Entropy loss, placing emphasis on accurate pixel-wise classification as the primary optimization objective.
-Table 1. Comparison with state-of-the-art parameter-efficient fine-tuning (PEFT) methods.
+Input images are uniformly resized to $1024\times 1024$ during both training and testing.
+The model is trained for 200 epochs using the Adam optimizer with a learning rate of $5\times 10^{-4}$ and a batch size of 2.
+For the segmentation task, we employ the standard Cross-Entropy loss, placing emphasis on accurate pixel-wise classification as the primary optimization objective.
+Table 1. Quantitative Comparison with State-of-the-Art Parameter-Efficient Fine-Tuning (PEFT) Methods
 Methods IoU Dice Param
@@
 Earth Adapter (2025) (Hu et al., 2025) 0.5149 0.5893 10.19M
-LoRA (2021) (Hu and Shen, 2022) 0.5202 0.5909 0.48M
-SAM Adapter (2023) (Wu et al., 2023b) 0.5321 0.6116 5.64M
+LoRA (2021) (Hu et al., 2022) 0.5202 0.5909 0.48M
+SAM Adapter (2023) (Wu et al., 2025) 0.5321 0.6116 5.64M
 Conv-LoRA (2024) (Zhong et al., 2024) 0.5342 0.6051 0.50M
 Ours 0.5520 0.6331 4.12M
-Figure 5. Visualization of the State-of-the-Art LULC Method.
+Figure 5.
+Visualization of the State-of-the-Art LULC Method.
+The results show that earlier methods do not work well on ultra-high spatial resolution data: (1) they have trouble outlining building boundaries, for example, nearby structures are often blurred together; (2) they miss small objects, such as residential swimming pools; and (3) they mix up similar classes, for instance confusing roads with parking lots.
+In contrast, our method gives clear improvements on all these points.
 4.4. Comparison With State-of-the-Art PEFT Methods
-To evaluate the effectiveness of the proposed adapter, we conducted extensive comparative experiments against several state-of-the-art parameter-efficient fine-tuning (PEFT) methods, including the SAM Adapter(Wu et al., 2023b), LoRA(Hu and Shen, 2022), Conv-LoRA(Zhong et al., 2024), and the Earth Adapter(Hu et al., 2025).
-These approaches represent diverse strategies for adapting large-scale models like SAM2, with different focuses on domain specificity and parameter efficiency, as summarized in Table 1.
-Limitations of LoRA and Conv-LoRA.
-Our experimental results show that both LoRA and Conv-LoRA achieve competitive performance with remarkably low parameter overhead.
-LoRA, which leverages low-rank decomposition for efficient adaptation, is widely adopted for lightweight fine-tuning tasks.
-However, due to its linear structure, it exhibits limited capability in modeling the spatial complexity of high-resolution remote sensing scenes, which often involve multi-scale objects, intricate semantics, and fine-grained textures.
-To address this limitation, Conv-LoRA enhances the original LoRA design by incorporating convolutional layers to strengthen local feature extraction.
-This modification improves its ability to model small objects and boundary details, leading to noticeable performance gains.
-Nonetheless, despite these improvements, Conv-LoRA still lacks the representational capacity to fully exploit the potential of high-resolution remote sensing imagery, particularly in capturing complex spatial-frequency interactions, leaving room for further enhancement.
-Limitations of Earth Adapter and SAM Adapter.
-The Earth Adapter, although specifically designed for remote sensing tasks, attempts to decouple frequency components without incorporating targeted spatial modeling mechanisms.
-This results in relatively high parameter consumption and unsatisfactory performance.
-The SAM Adapter, originally developed for medical image segmentation, shows moderate adaptability when transferred to remote sensing domains.
-However, due to the increased complexity and resolution of remote sensing data, it struggles to effectively model fine-grained structures and semantic variations.
-Effectiveness of the Proposed FreqWeaver Adapter.
-In contrast, our proposed FreqWeaver Adapter introduces a frequency-aware design tailored for remote sensing applications.
-By explicitly decoupling low- and high-frequency components and applying targeted processing strategies to each, FreqWeaver effectively suppresses low-frequency artifacts while preserving high-frequency structural details.
-This leads to improved segmentation accuracy and better adaptation to the SAM2 backbone.
-Importantly, FreqWeaver achieves this superior performance with fewer parameters than both the Earth Adapter and the SAM Adapter, demonstrating its strong domain-specific effectiveness and efficiency for high-resolution remote sensing tasks.
-We also present visualizations to validate the effectiveness of our method, as shown in 3.
-Table 2. Comparison with state-of-the-art Land Use and Land Cover (LULC) methods.
+To evaluate the effectiveness of the proposed adapter, we conduct extensive comparative experiments against several state-of-the-art parameter-efficient fine-tuning (PEFT) methods, including the SAM Adapter (Wu et al., 2025), LoRA (Hu et al., 2022), Conv-LoRA (Zhong et al., 2024), and the Earth Adapter (Hu et al., 2025).
+All implementations are based on SAM2, ensuring a consistent backbone for fair comparison.
+These approaches represent diverse strategies for adapting SAM2, with different focuses on domain specificity and parameter efficiency.
+The comparison results are shown in Table 1.
+Limitations of LoRA-Based Methods.
+Our experimental results show that both LoRA and Conv-LoRA achieve competitive performance with very low parameter overhead.
+LoRA leverages low-rank decomposition for efficient adaptation and is widely used in lightweight fine-tuning tasks.
+However, due to its linear structure, it has limited ability to capture the spatial complexity of high-resolution remote sensing scenes, which often contain multi-scale objects, intricate semantics, and fine-grained textures.
+To mitigate this limitation, Conv-LoRA extends the original LoRA design by incorporating convolutional layers to enhance local feature extraction.
+This modification improves its capacity to model small objects and boundary details, resulting in measurable performance gains.
+Nevertheless, despite these improvements, Conv-LoRA still falls short in fully exploiting the potential of high-resolution remote sensing imagery, particularly in capturing multi-scale spatial structures and fine-grained textures, which indicates the need for further refinement.
+Limitations of Adapter-Based Methods.
+The SAM Adapter is originally developed for medical image segmentation and shows a certain level of adaptability when transferred to remote sensing tasks.
+However, medical images usually contain relatively simple structures with clear boundaries, while remote sensing data have higher resolution, more complex scene variations, and richer semantic categories.
+As a result, the SAM Adapter has limited ability to model fine-grained structures and semantic differences, and its performance improvement remains modest.
+In contrast, the Earth Adapter is specifically designed for remote sensing tasks, but its attempt to decouple frequency components lacks targeted high- and low-frequency modeling mechanisms.
+This limitation prevents it from fully capturing complex spatial structures and texture features, leading to higher parameter overhead and unsatisfactory overall performance.
+Effectiveness of FreqWeaver Adapter.
+Our FreqWeaver Adapter introduces a frequency-aware design for remote sensing applications.
+It explicitly decouples low- and high-frequency components and applies targeted processing to each, which suppresses low-frequency artifacts while preserving high-frequency structural details.
+This improves segmentation accuracy and allows better adaptation to the SAM2 backbone.
+Compared with the best-performing Conv-LoRA, FreqWeaver increases IoU by 1.78% and Dice by 2.8%, showing clear gains in both overall segmentation accuracy and fine-grained detail modeling.
+In addition, FreqWeaver achieves this higher performance with fewer parameters than both the Earth Adapter and the SAM Adapter, demonstrating strong domain-specific effectiveness and efficiency for high-resolution remote sensing tasks.
+We also provide visualizations to further validate its effectiveness, as shown in Fig.
+4.
+Table 2. Quantitative Comparison with state-of-the-art Land Use and Land Cover (LULC) methods.
 Methods IoU Dice
@@
 UNetFormer(2022)(Wang et al., 2022) 0.5236 0.6201
-CMTFNet(2023)(Wu et al., 2023a) 0.5428 0.6369
+CMTFNet(2023)(Wu et al., 2023) 0.5428 0.6369
 RS3Mamba(2024)(Ma et al., 2024b) 0.5092 0.6068
@@
 Ours 0.5778 0.6545
-4.5. Comparison with State-of-the-Art Land Cover Classification Methods
-To evaluate the effectiveness of our proposed approach, we conducted extensive comparative experiments against several representative state-of-the-art Land Cover and Land Use Classification (LULC) methods.
-Specifically, we selected five recent and competitive baselines from the literature (2021–2024)—ABCNet (Li et al., 2021), UNetFormer (Wang et al., 2022), CMTFNet (Wu et al., 2023a), RS3Mamba (Ma et al., 2024b), and SAM-RS (Ma et al., 2024a) chosen for their strong benchmark performance and representation of diverse architectural paradigms, including CNN-based, Transformer-based, and Mamba-based models.
-The quantitative comparison results are reported in Table 2, while qualitative visualization examples are provided in Fig.
+4.5. Comparison with State-of-the-Art LULC Methods
+To evaluate the effectiveness of our proposed approach, we conduct extensive comparative experiments against several representative state-of-the-art Land Cover and Land Use Classification (LULC) methods.
+Specifically, we select five recent and competitive baselines from the literature (2021–2024)—ABCNet (Li et al., 2021), UNetFormer (Wang et al., 2022), CMTFNet (Wu et al., 2023), RS3Mamba (Ma et al., 2024b), and SAM-RS (Ma et al., 2024a)—chosen for their strong benchmark performance and representation of diverse architectural paradigms, including CNN-based, Transformer-based, and Mamba-based models.
+The quantitative comparison results are shown in Table 2, while qualitative visualization examples are presented in Fig.
 5 to further illustrate the advantages of our method.
@@
 In comparison to the strongest baseline, SAM-RS, our model achieves an improvement of 3.44% in IoU and 2.03% in Dice score, underscoring its superior performance in LULC tasks.
-Despite the reasonable segmentation accuracy exhibited by prior methods, visual analysis reveals several limitations: (1) Misclassification remains prevalent in complex scenes, particularly those with high intra-class variance or subtle inter-class distinctions (e.g., differentiating various types of vegetation or urban structures); (2) Object boundaries are often rendered blurry and imprecise, a critical deficiency for ultra-high-resolution segmentation tasks that demand sharp delineations; (3) Models generally perform well on spectrally distinct classes like grass and trees, but falter with categories exhibiting complex geometries or semantic overlaps, such as distinguishing between parking lots and buildings, or between roads and barren land.
+Despite the reasonable segmentation accuracy exhibited by prior methods, visual analysis reveals several limitations: (1) Misclassification remains prevalent in complex scenes, particularly those with high intra-class variance or subtle inter-class distinctions (e.g., differentiating various types of vegetation or urban structures); (2) Object boundaries are often blurry and imprecise, a critical deficiency for ultra-high-resolution segmentation tasks that demand sharp delineations; (3) Models generally perform well on spectrally distinct classes like grass and trees, but falter on categories with complex geometries or semantic overlaps, such as distinguishing between parking lots and buildings, or between roads and barren land.
 Moreover, small object detection, such as identifying pools in residential areas, is often overlooked by existing approaches.
 Effectiveness of Our Method.
-To the best of our knowledge, our model is the first to adapt the Segment Anything v2 (SAM2) architecture for the task of LULC.
-By leveraging the rich prior knowledge encoded in this foundation model, we obtain robust category- and pixel-level representations.
-Building upon this foundation, the proposed FreqWeaver Adapter enhances the semantic granularity of feature modeling—by explicitly capturing and integrating multi-frequency information crucial for detail preservation—thereby enabling significantly improved delineation of object boundaries and detection of small structures.
-Notably, the enhanced feature representation from the FreqWeaver Adapter allows our model to effectively exploit contextual cues, such as the presence or absence of vehicles, to better distinguish between semantically similar regions like parking lots and roads.
+We adapt the SAM2 foundation model for the LULC task, leveraging its rich prior knowledge to obtain robust category- and pixel-level representations.
+Building upon this foundation, our FreqWeaver Adapter enhances the semantic granularity of feature modeling by explicitly capturing and integrating multi-frequency information crucial for detail preservation, thereby enabling significantly improved delineation of object boundaries and detection of small structures.
+Notably, the enhanced feature representation from the FreqWeaver Adapter allows our model to effectively exploit contextual cues to better distinguish between semantically similar regions such as parking lots and roads.
 These capabilities demonstrate the strengths of our architecture in ultra-high-resolution LULC tasks.
-Additionally, the integration of a semi-supervised learning strategy further improves the robustness and generalization of our model under limited annotation scenarios.
+In addition, the integration of a semi-supervised learning strategy further improves the robustness and generalization of our model under limited annotation scenarios.
 Table 3. Classification performance on ultra-high resolution Downstream Applications.
@@
 Kappa Coefficient 81.38%
-Figure 6. Visualization of Land Cover Classification.
-4.6. Exploration of Downstream Applications in Map Analysis
-To evaluate the practicality of our approach in real-world scenarios, we applied the proposed method to downstream ultra-high-resolution map analysis tasks, as summarized in Table 3.
+Figure 6.
+Visualization of Land Cover Classification on the Baltimore City Map.
+The results show that our method achieves reliable land cover classification across the entire city at ultra-high spatial resolution, clearly revealing the overall urban structure.
+4.6. Downstream Applications in Ultra-High Spatial Resolution Map Analysis
+To evaluate the practicality of our approach in real-world scenarios, we apply the proposed method to downstream ultra-high-resolution map analysis tasks, as summarized in Table 3.
 Despite being trained on a relatively small dataset, our method exhibits strong generalization capabilities, achieving an Overall Accuracy (OA) of 84.55% and a Kappa coefficient of 81.38%.
 These metrics reflect not only the correctness of predictions but also the statistical reliability of classification performance across diverse land cover categories.
-A detailed breakdown of class-wise performance is provided in Table 3.
-Notably, high User’s Accuracy was achieved for categories such as Building (93.71%), Water (94.92%), and Grass/Shrub (91.80%), indicating precise identification of spatial regions with distinct spectral or structural characteristics.
-At the same time, balanced Producer’s Accuracy was maintained, particularly for Tree Canopy (98.98%) and Road (89.94%), demonstrating strong recall and comprehensive spatial coverage.
+A detailed breakdown of class-wise performance is shown in Table 3.
+Notably, our model achieves high User’s Accuracy for categories such as Building (93.71%), Water (94.92%), and Grass/Shrub (91.80%), indicating precise identification of spatial regions with distinct spectral or structural characteristics.
+At the same time, it maintains balanced Producer’s Accuracy, particularly for Tree Canopy (98.98%) and Road (89.94%), demonstrating strong recall and comprehensive spatial coverage.
 Nevertheless, certain challenges persist in distinguishing semantically overlapping or visually ambiguous regions, such as Tree Canopy vs.
@@
 Road, which are frequently misclassified due to spatial proximity or similar textures.
-To further validate the effectiveness of our method, qualitative visualizations were performed on real-world map segments, as illustrated in Fig.
+To further validate the effectiveness of our method, we perform qualitative visualizations on real-world map segments, as illustrated in Fig.
 6.
 The results show that our approach successfully preserves object boundaries, captures fine-grained structures, and adapts effectively to spatial heterogeneity, confirming its suitability for downstream high-resolution map analysis tasks.
-Table 4. Quantitative results of the ablation study.
-FW UATS IoU Dice
-0.5239 0.6006
-0.5520 0.6331
-0.5778 0.6545
 4.7. Ablation Studies
-To evaluate the contribution of each component within the proposed framework, we conducted extensive ablation studies, as shown in Table 4.
-In each experiment, specific modules—FreqWeaver (FW) and the Uncertainty-Aware Teacher Student Model (UATS)—were selectively enabled or disabled to analyze their individual impact on segmentation performance.
+To evaluate the contribution of each component within the proposed framework, we conduct extensive ablation studies, as shown in Table 4.
+In each experiment, specific modules—FreqWeaver (FW) and the Uncertainty-Aware Teacher–Student Framework (UATS)—are selectively enabled or disabled to analyze their individual impact on segmentation performance.
 Effectiveness of FreqWeaver.
-We introduce FreqWeaver (FW) to decouple the frequency and temporal components within SAM2, enabling a parameter-efficient mechanism to capture complex patterns in remote sensing imagery.
+We introduce FreqWeaver (FW) to decouple the frequency components and apply specific modeling within SAM2, enabling a parameter-efficient mechanism to capture complex patterns in remote sensing imagery.
 Experimental results show that, compared to full fine-tuning, our method achieves a 2.81% improvement in IoU and a 3.25% improvement in Dice score, while updating only 5.96% of the total parameters.
@@
 These results demonstrate the importance of uncertainty modeling in enhancing model robustness and generalization across diverse remote sensing scenes.
+Table 4. Quantitative results of the ablation study.
+FW UATS IoU Dice
+0.5239 0.6006
+0.5520 0.6331
+0.5778 0.6545
 Table 5. Quantitative results of the Adaptation ablation study.
@@
 However, directly applying this strategy to Segment Anything 2 often results in suboptimal fine-tuning performance due to its deeper and hierarchically structured design.
-To better accommodate these architectural differences, we conducted a series of empirical studies to investigate more effective adapter integration strategies.
-The results are presented in Table 5.
+To better accommodate these architectural differences, we conduct a series of empirical studies to investigate more effective adapter integration strategies.
+The results are shown in Table 5.
 Specifically, we hypothesize that fine-tuning a foundation model to adapt to downstream tasks such as remote sensing segmentation primarily requires the adjustment of deeper semantic layers.
-Therefore, we initially inserted adapters into all four stages and gradually moved toward inserting them only in the deeper stages.
-Experimental results reveal that adding adapters to early transformer stages not only fails to improve domain adaptation, but can even hinder the model’s ability to specialize for remote sensing data.
+Therefore, we initially insert adapters into all four stages and gradually move toward inserting them only in the deeper stages.
+Experimental results reveal that adding adapters to early transformer stages not only fails to improve domain adaptation but can even hinder the model’s ability to specialize for remote sensing data.
 Conversely, limiting adapter placement to only the final stage also leads to underfitting, as the model lacks sufficient capacity for structural adaptation.
 We observe that inserting adapters into Stage 3 and Stage 4 yields the best performance, striking a balance between semantic depth and structural flexibility.
-While our experiments primarily report results using the proposed FreqWeaver Adapter, we also validated this stage-wise trend using other adapter designs, including the SAM Adapter and Earth Adapter, and found consistent behavior across all cases.
+While our experiments primarily report results using the proposed FreqWeaver Adapter, we also validate this stage-wise trend using other adapter designs, including the SAM Adapter and Earth Adapter, and find consistent behavior across all cases.
 5. Conclusion
-In this work, we proposed a parameter-efficient semi-supervised segmentation framework for high-resolution remote sensing imagery.
-By integrating a domain-adaptive FreqWeaver Adapter with the SAM2 backbone and employing an uncertainty-aware teacher student strategy, our method addresses the challenges of limited annotations, scale variation, and structural consistency.
-Experiments show that our approach achieves strong segmentation performance with minimal parameter overhead, outperforming existing PEFT and high-resolution segmentation methods.
-These results highlight the potential of combining spectral decomposition and semi-supervised learning for effective foundation model adaptation in remote sensing.
+In this work, we present Baltimore Atlas, a framework that combines the FreqWeaver Adapter and the Uncertainty-Aware Teacher Student Framework for ultra-high spatial resolution land cover classification.
+The FreqWeaver Adapter separates image information into high- and low-frequency components, modeling each to capture both global and local details, and adapts SAM2 to UHSR data with minimal parameters.
+The Uncertainty-Aware Teacher–Student Framework uses high-confidence pseudo-labels to learn more reliable features from unlabeled data.
+Combined, these components support fine-grained UHSR land cover classification and reduce the need for large labeled datasets.
+Experiments on the Baltimore Atlas dataset show that our FreqWeaver Adapter outperforms leading parameter-efficient tuning methods, and the full framework achieves better results than current state-of-the-art land cover classification approaches.
 6. Discussion